### v1

# GOAL: A Generalist Combinatorial Optimization Agent Learner

Abstract
Machine Learning-based heuristics have recently shown impressive performance in solving a variety of hard combinatorial optimization problems (COPs). However they generally rely on a separate neural model, specialized and trained for each single problem. Any variation of a problem requires adjustment of its model and re-training from scratch. In this paper, we propose GOAL (for Generalist combinatorial Optimization Agent Learning), a generalist model capable of efficiently solving multiple COPs and which can be fine-tuned to solve new COPs. GOAL consists of a single backbone plus light-weight problem-specific adapters, mostly for input and output processing. The backbone is based on a new form of mixed-attention blocks which allows to handle problems defined on graphs with arbitrary combinations of node, edge and instance-level features. Additionally, problems which involve heterogeneous nodes or edges, such as in multi-partite graphs, are handled through a novel multi-type transformer architecture, where the attention blocks are duplicated to attend only the relevant combination of types while relying on the same shared parameters. We train GOAL on a set of routing, scheduling and classic graph problems and show that it is only slightly inferior to the specialized baselines while being the first multi-task model that solves a variety of COPs. Finally we showcase the strong transfer learning capacity of GOAL by fine-tuning or learning the adapters for new problems, with only few shots and little data.
1 Introduction
Combinatorial Optimization problems are extensively studied in the Operations Research community, because they are both challenging in terms of theoretical complexity and central in many applications such as transportation, logistics and finance. Traditional methods include exact solvers (typically based on integer programming [32]) and approximation algorithms [55] which provide optimality or approximation guarantees, respectively, but do not scale; and heuristic algorithms [9] which produce good-quality solutions faster but without guarantees. Effective heuristics heavily rely on expert knowledge and are generally problem-specific. A more recent and promising family of methods for CO is based on machine learning, either to learn the hyper-parameters of classic CO solvers, or to directly learn heuristics in an end-to-end fashion [6, 11]. In particular constructive neural heuristics have been successful for solving a variety of CO problems, including routing [5, 42, 30, 33], scheduling [58, 34] and graph problems [27, 45]. However, in most cases, a separate model is specialized and trained for each problem, and sometimes even for each distribution of instances within a problem, especially for varying instance sizes. While much effort has been devoted to improving the generalization to different instance distributions [41, 59, 49, 16], generalization to different problems is less studied. Indeed, although there are strong neural heuristics for many CO problems, any variation of these problems requires adjustment of their model and re-training from scratch. While it is true that some problem-specific model tailoring is always required, since the structure of problem instances and solutions are very diverse, many problems are related and it would be interesting to exploit their commonalities. In this paper, we seek to answer the following question: Can one model be pretrained in order to be easily specialized to solve a variety of CO problems?
The idea of pretraining a unified model on some pretext task so that it can be efficiently fine-tuned to several downstream tasks has been extremely successful in the language and vision domains [15, 10, 46]. In these domains, a self-supervised pretraining task exploits the intrinsic structure and regularities of images and text to learn generic representations. In contrast, in CO, while most problem instances can be modeled as graphs, there is no generic regularity to leverage by a self-supervised task. Graphs, unlike images or text, are a “poor” modality, with little redundancy or intrinsic semantics.
In this paper, we propose GOAL (for Generalist combinatorial Optimization Agent Learning), a generalist model capable of efficiently solving a variety of CO problems and which can be fine-tuned to solve new related problems. GOAL consists of a single backbone and light-weight problem-specific adapters for input and output processing. To handle the large discrepancy between CO problems, we propose two complementary strategies. First, the backbone is based on a new form of mixed-attention blocks which allows to handle problems defined on graphs with arbitrary combinations of node, edge and instance-level features. Additionally, problems which involve heterogeneous nodes or edges, which typically arise with multipartite graphs, are handled through a novel multi-type transformer architecture, where the backbone mixed-attention blocks are duplicated to attend each relevant combination of types while relying on the same shared parameters. GOAL is able to handle any CO problem whose instances can be represented as a graph and where feasible solutions can be constructed iteratively, as a sequence of node selections. We find that this pattern is generic enough to cover a wide range of problems. We train GOAL in a multi-task setting, by imitation of expert trajectories of eight classic CO problems, spanning multiple domains: Routing problems: the Asymmetric Traveling Salesman Problem (ATSP), the Capacitated Vehicle Routing Problem (CVRP), the CVRP with Time Windows (CVRPTW), the Orienteering Problem (OP); Scheduling problems: the Job Shop Scheduling Problem (JSSP) and the Unrelated Machine Scheduling Problem (UMSP); Packing problems: the Knapsack Problem (KP); and Graph problems: the Minimum Vertex Covering Problem (MVC). The expert trajectories are based on good-quality solutions, easily generated by traditional problem-specific solvers for instances of medium size (generally 100 nodes).
Experimentally we first show that GOAL performs very well on all the training tasks, making it the first model that solves such a variety of CO problems. Second, we illustrate GOAL’s strong zero (or few) shot generalization performance to various distributions shifts from the training tasks’ instances. Third, we fine-tune GOAL on four new problems and demonstrate that it achieves a good performance quickly using little data. We also discuss scenarios where fine-tuning is less effective, hinting at the required conditions for successful transfer; namely that the new task is close enough to at least one of the training tasks to share a similar instance structure, while its objective and constraints can be completely different. We further provide an ablation study to evaluate the effectiveness of our multi-type transformer architecture and mixed-attention blocks. Notably our architecture gives state-of-the-art results on 6 of our 8 training tasks when used in a single-task version of GOAL. In summary, our contributions are the following:
  • We introduce GOAL: a generic and flexible model for end-to-end combinatorial optimization. Two key ingredients of GOAL are the simple yet effective form of mixed-attention blocks which enables handling problems represented by graphs with any combination of node, edge and instance-level features and its multi-type transformer architecture which allows to seamlessly tackle problems with multipartite graph structures.
  • We demonstrate that it is possible to learn a single backbone model to effectively solve a variety of CO tasks, spanning routing, scheduling, packing and graph problems.
  • We showcase the potential of such a multi-task pretrained model to be efficiently adapted to handle new CO tasks and provide empirical insights on how closely related the new tasks should be to the training tasks to ensure successful transfer.
2 Related Works
End-to-end heuristic learning for CO
Our work falls into the category of end-to-end constructive heuristics for combinatorial optimization. Constructive neural heuristics have been successfully applied to a wide range of COPs, including many variants of routing problems [30, 57, 28, 31, 18, 52], scheduling problems [58, 34], as well as classic COPs such as the Knapsack Problem [33] and the Maximum Independent Set [27, 52]. Another successful category of neural CO methods are improvement heuristics, which focus on iteratively enhancing a given complete solution, typically by using a neural network to guide the selection of local search operators [13, 22, 40, 56]. The seminal work of [27] introduces graph embeddings which do work for a variety of CO problems but with limited performance. Current state-of-the-art models are tailored for certain problem structures (e.g. Euclidian routing problems, where nodes are represented by their coordinates plus some extra node or graph features such as demand or vehicle capacity), and successfully handled by transformer-based architectures [30]. For non-euclidian routing problems, [34] and [16] propose adapted architectures to deal with the distance matrix, but they are specialized for such problems and do not apply to those with node-only features. In GOAL, we propose a novel mixed-attention block which can flexibly deal with both node and edge features, incurring no overhead when one or the other is missing.
Generalization of neural heuristics
While the first seminal models for neural CO had a dramatic drop in performance when tested on out-of-training distribution instances [26, 41], several recent approaches show a greatly improved generalization performance, especially to larger instances [18, 35, 49, 43], to distribution shifts of other instance parameters [8], or both [41, 8, 59, 16]. Concurrently with our work, some papers have started to explore the cross-problem generalization for the family of routing problems. Specifically, [36] propose to pretrain the AM model [30] on the TSP and to fine-tune either the entire model or some light-weight modules on the Price Collecting TSP and OP. Additionally, [60] and [37] explore generalization to different variants of the CVRP. They build on the idea of zero-shot compositionality [48] by training a model on CVRPs with some combinations of constraints from a predefined set (e.g. time windows and duration limit constraints) and show that the model generalizes to some unseen combinations of the same constraints. Although these approaches tackle multiple problems, it is important to note that the structures of the considered Euclidian routing problems are very similar and indeed many of those problems were solved by the seminal Attention Model of [30] with little adaptations of the base model. Our paper goes further and considers multi-task training and transfer learning of a new model on a variety of COPs with different structure.
Multi-task training for sequential decision making
Multi-task (pre-)training has recently shown promising results in related fields. For example, [25] proposed a graph neural network processor capable of learning to execute a variety of algorithms, such as sorting and searching. The paper focuses only on polynomial problems and learns to imitate the execution of classical algorithms. In contrast, we address NP-hard problems and learn by imitation of the solution provided by (potentially exponential) oracle algorithms. Recent approaches in control rely on sequence modeling of expert trajectories from multiple tasks [50, 51]. Although they treat various tasks, with potentially diverse state and action spaces, these works still consider image observations and exploit this structure. [47] proposes GATO, a general-purpose agent obtained by training a large transformer-based model on a hundreds of tasks spanning vision, language and control. It demonstrates the potential of a single network (albeit with 1.2B parameters) to effectively tackle a wide variety of tasks.
3 Multi-Task Learning for Combinatorial Optimization
Refer to caption Refer to caption
Figure 1: Left: Architecture of GOAL for single-type problems; the green components are task-specific, while the others belong to the backbone. Right: Architecture of GOAL for multi-type problems (with here 2 types of nodes T1,T2); “self” and “cross” are identical multi-head mixed attention blocks (same parameters for all types and tasks); “FF” for the two types are also identical Feed Forward blocks; for the edges, all the type combinations are shown (here T11,T12,T21,T22) but each task can define its own. The same set of parameters is used for the two architectures.
3.1 Constructive CO problem solving
We assume that a CO problem instance is given by a pair of a finite, non-empty set $X$ of feasible solutions, and a real-valued objective function $f$ whose domain contains $X$. Solving the instance consists in finding an optimal solution, i.e. an element of $\arg\min_{x\in X}f(x)$. Constructive approaches assume that $f$ is defined not only over $X$ but over a set of partial solutions, which contains $X$ and is amenable to a step-wise construction procedure: each partial solution can be obtained by a sequence of construction steps. Thus, solving the problem instance can be formulated as a sequential decision procedure which iteratively selects the construction steps so as to reach not only a feasible but an optimal solution. That in turn can be modeled as a Markov Decision Problem (MDP) associated with the problem instance.
The action space of this MDP is naturally defined as the (discrete) set of construction steps, but there are many ways to choose its state space. For the very broad class of so called “tail-recursive” CO problems, [16] shows that an effective choice is the space of all possible instances of the problem. Indeed, the construction of a solution can be modeled as a trajectory in the space of problem instances, following the tail sub-problems [7] obtained after each construction step. All the classical standard NP-hard CO problems, such as vehicle routing problems, machine scheduling, various graph problems etc. are tail-recursive, sometimes modulo a slight reformulation (e.g. TSP is not per se tail-recursive, but path-TSP is). In the sequel, we deal exclusively with tail-recursive CO problems and their construction MDPs, called BQ-MDP in [16].
3.2 Multi-Task CO Problems
Now, given a class $\mathcal{T}$ of tail-recursive CO problems called tasks, let $\boldsymbol{\Omega}_{t}$ for each task $t{\in}\mathcal{T}$ denote the parametric space used to represent its instances. Consider the new (multi-task) CO problem defined as the union of the tasks in $\mathcal{T}$ (each being a set of instances). Its instances can be represented in the disjoint union $\biguplus_{t\in\mathcal{T}}\boldsymbol{\Omega}_{t}$ consisting of pairs $(t,\omega)$ of a task identifier $t{\in}\mathcal{T}$ and a point $\omega{\in}\boldsymbol{\Omega}_{t}$ in its corresponding parameter space. It is easy to show that the multi-task problem is tail-recursive whenever each of its components is. And its BQ-MDP is then simply the disjoint union of the BQ-MDPs of its components.
In this paper, we study alternatives for the multi-task policy space, beyond the trivial one where no parameter is shared between the policies of the different tasks (Cartesian product of the policy spaces of the individual tasks). The underlying assumption is that if the tasks are sufficiently similar, learning one task may elicit implicit skills useful to other tasks, which are then leveraged by parameter sharing. In the extreme, some tasks may require little training data and/or just a few shot fine-tuning to obtain acceptable performance, leveraging their similarity to more thoroughly trained tasks. Since each task has its own parametric space used to represent its instances, some part of any multi-task policy has to be task specific, and trained by processing at least a few instances of each task: true zero-shot generalization is not possible. The situation is close to multi-lingual Natural Language Processing, where the vocabulary of each language needs to be seen at least once at training. And we borrow the solution used in that context, namely adapters [4], to capture the task-specific parts of the policy.
3.3 Task Representation
To leverage operational similarity between tasks, we need them to have at least representational similarity. Thus, we assume that each instance of size $N$ of a task (CO problem) is defined by a set of $N$ abstract “nodes”, together with features which can be attached at node, edge (node pair) or instance level. Instance features can be replicated and appended to the features of each node, hence can be ignored without loss of generality. Since the construction steps (actions) are returned by the task-specific output adapters (Sec 3.4), there are no restrictions on their format. For simplicity and to cover the problems used in our experiments, we assume in the sequel that each construction step consists of the selection of a node, possibly with an option among a fixed number $K$ of options (dependent on the task: in most cases, $K{=}1$ and the action simply selects a node).
Hence, the GOAL model takes three inputs $t,x,\bar{x}$, where $t$ is a task identifier, and $(x,\bar{x}){\in}\boldsymbol{\Omega}_{t}$ is a parameter representing an instance of task $t$, where $x{\in}\mathbb{R}^{N{\times}F_{t}}$ (resp. $\bar{x}{\in}\mathbb{R}^{N{\times}N{\times}\bar{F}_{t}}$) represents the feature vectors of dimension $F_{t}$ of the $N$ nodes (resp. $\bar{F}_{t}$ of the $N{\times}N$ edges). While some tasks may have no edge feature ($\bar{F}_{t}{=}0$), we always assume that at least one node feature exists, which distinguishes the nodes by a randomly assigned value between 00 and $1$, acting as node identifier, so we always have $F_{t}{\geq}1$. Both $F_{t}$ and $\bar{F}_{t}$, as well as the meaning of the corresponding features, depend on task $t$. The output of the model is a matrix $y{\in}\mathbb{R}^{N{\times}K_{t}}$ representing the score of each candidate action, i.e. the pair of a node and an option for selection. As usual, it is passed to a softmax (over the whole matrix) to obtain a probability distribution. Wherever appropriate, task-dependent masking can be used to ensure that probabilities are null for the actions disallowed by a task.
For example, for the path-TSP task, the nodes are the customers to visit plus the origin and destination of the vehicle. There are two proper node features, each holding a binary value indicating whether the node is the origin or the destination, respectively (thus $F_{t}{=}3$ counting node identifier). And there is a single edge feature ($\bar{F}_{t}{=}1$) holding the edge cost (distance, time, energy, etc.). The action is a simple node selection ($K_{t}{=}1$), with origin and destination nodes being masked. In a BQ-MDP transition, the selected node becomes the new origin, and the previous origin is discarded ($N$ is decreased by $1$).
3.4 Architecture of the Model
The architecture of our model generalizes that of BQ-NCO [16], and is described in Fig. 1 left. It essentially consists of a sequence of layers whose parameters are shared by all the tasks (backbone) together with task-specific adapter modules at the input and output of the model [^1] Additional adapters inside layers can also be introduced, but this paper considers only input-output adapters.. The input-output adapter modules for task $t$ are specific to the dimensions $F_{t},\bar{F}_{t},K_{t}$ and the meaning of the corresponding features. No positional encoding is used (all relational information is assumed to be captured by $\bar{x}$). The layers themselves, on the other hand, are only aware of the shared embedding dimensions $D,\bar{D}$. They are self-attention transformer layers [54] where vanilla attention is replaced by multi-head mixed attention, similar in intent to [34] but with a different design.
Our mixed attention block, in the general case where $M$ nodes attend on $N$ nodes ($M{=}N$ in self-attention), is an instance of generalized attention [2]. It takes the following inputs: node representations as vectors for $N$ queries, $M$ keys, and $M$ values, with an optional $M{\times}N$ log-binary mask $\mathcal{M}$, as in vanilla attention, plus optional edge representations as $M{\times}N$ “edge” vectors. In each head $h$, each query, key, value vector is first projected, as in vanilla attention, into a common space of dimension $d$, yielding for each node $n{=}1{:}N$ and $m{=}1{:}M$, vectors:
$Q_{n}^{(h)}=Q_{n}\mathbf{W}_{Q}^{(h)},\hskip 14.22636ptK_{m}^{(h)}=K_{m}\mathbf{W}_{K}^{(h)},\hskip 14.22636ptV_{m}^{(h)}=V_{m}\mathbf{W}_{V}^{(h)}.$
The edge vectors are also projected into the same space of dimension $d$, but twice so: once as a key and once as a query, resulting in vectors:
$Q^{\prime(h)}_{mn}=E_{mn}\mathbf{W}^{\prime(h)}_{Q},\hskip 14.22636ptK^{\prime(h)}_{mn}=E_{mn}\mathbf{W}^{\prime(h)}_{K}.$
Vanilla attention then builds an $M{\times}N$ score matrix $S_{mn}^{(h)}{=}Q_{n}^{(h)}K_{m}^{(h)\top}$, from which the attention coefficients are computed (using softmax and involving the mask). We simply augment the factors of the scalar product defining the score matrix with the edge projections, and keep the rest unchanged, yielding new representations $r$ of the $N$ nodes:
$\begin{split}r=\sum_{h}\textrm{softmax}_{\textrm{col}}(S^{(h)}{+}\mathcal{M})^{\top}V^{(h)}\mathbf{W}_{O}^{(h)\top}\hskip 2.84544pt\\ \text{ with }\hskip 2.84544ptS^{(h)}_{mn}{=}(Q_{n}^{(h)}{+}Q^{\prime(h)}_{mn})(K_{m}^{(h)}{+}K^{\prime(h)}_{mn})^{\top}.\end{split}$ (1)
In GOAL, all the node inputs are of dimension $D$, the edge inputs of dimension $\bar{D}$, and they are all projected into a space of dimension $d$ equal to $D$ divided by the number of heads. Furthermore, no masking is used in the attention blocks.
One well-known limitation of this model is the quadratic complexity bottleneck of attention matrices, which are of size $N^{2}$. In practice though, for many tasks, a simple, task-specific, heuristic can deal with that problem. Essentially, instead of presenting all the $N$ nodes of the input instance to the model, the heuristic selects a priori a subset of the $n$ most promising ones. Thus, the attention matrices never exceed size $n^{2}$, even when $N^{2}$ would be unmanageable. The performance impact is often negligible, and even sometimes improving (maybe due to the dilution of attention coefficient when $N$ is large). For example, with TSP, only the origin, destination and 100-nearest neighbors to the origin are presented.
3.5 Multi-Type Transformer Architecture
Many tasks (CO problems) involve different types of nodes. For example, scheduling problems often involve “operation” nodes (the things to schedule) and “machine” nodes (which constrain feasible schedules). For GOAL, one solution is to simply treat them all indiscriminately in the backbone. The adapters, being problem specific, are of course free to treat them differently according to their types. This is sometimes unsatisfactory though, since it forces the adapter to provide uniform embeddings for all the edges, though they are often heterogeneous in the input: in the scheduling example, edge features may be precedence constraints between operations, or processing times between machines and operations, and no information at all between machines. Instead of packing all these different types of relations into a single big tensor, with padding values wherever a feature is not defined for a given edge, we propose a multi-type architecture where the adapters output separate embeddings per type (for nodes) or compatible pairs of types (for edges). Each layer of the model, instead of being a monolithic self-attention block operating on all the nodes, becomes a task-specific combination of both cross- and self-attention blocks, as described by (1), operating on different type-compatible subsets of nodes and edges. This is illustrated in Fig. 1 right. Therefore, in our scheduling example, the model will contain separate self-attention blocks, for operation nodes and machine nodes respectively, and cross-attention blocks, for operation-machine edges and vice-versa. Note that even when edges are not oriented, it is useful to have two cross-attention blocks with transposed matrices as edge embeddings: the projections in (1) are then used to interpret differently the two directions of the edges in each attention block. Finally, observe that while the architecture configuration is now task-specific, the model parameters remain the same for all the possible configurations. This is precisely what enables us to train a unique model to solve various single and multi-type problems.
4 Experiments
Our experiments seek to answer the following questions: Does the proposed GOAL approach enable learning a single backbone to effectively solve a variety of CO problems? (Sec 4.1); Can we use it as a pretrained model and efficiently fine-tune it to solve new unseen CO problems? (Sec 4.2); and finally, to what extent do the introduced mixed-attention blocks and multi-type architecture contribute to the results? (Sec 4.3). We start by describing the multi-task training of GOAL.
Training tasks.
We train GOAL on eight classic and varied CO problems. For each problem, we generate a training dataset of 1 million random instances and solve them using traditional solvers. Specifically, we consider the following problems and oracle solvers: the Asymmetric Traveling Salesman Problem (ATSP) with LKH [19], the Capacitated Vehicle Routing Problem (CVRP) with LKH [19], the Capacitated Vehicle Routing Problem with Time Windows (CVRPTW) with HGS [19] the Orienteering Problem (OP) with A4OP [29], the Job Shop Scheduling Problem (JSSP) with ORTools [44], the Unrelated Machine Scheduling Problem (UMSP) with HiGHS [24], the Knapsack Problem (KP) with ORTools [44] and the Minimum Vertex Covering Problem (MVC) with FastWVC [1]. The training instances are of size 100 nodes, or 10$\times$10 for JSSP and 10$\times$100 for UMSP (resulting in both cases in 100 operation nodes and 10 machine nodes). For the sake of space, the description of the problems and associated data generation processes is provided in Appendix A.
Training process.
GOAL is trained by alternating batches from each of the training tasks (round robin), using a single-class cross entropy loss for ATSP, CVRP, CVRPTW, OP, JSSP and UMSP and a multi-class cross entropy loss for KP and MVC. Similarly to [16], we create a batch by sampling suffix subproblems for each instance. We use AdamW optimizer [38], with an initial learning rate of 0.0005 and a decay rate every 10th epochs. We trained the model on 8 Nvidia A100 GPU servers, with batch size of 128 for 7 days, allowing the completion of approximately 500 epochs. The best model was selected based on the average performance on a validation set of 512 instances per problem. The model hyperparameters are described in Appendix C.
Baselines.
We consider the following constructive neural models as baselines: AM [30], MDAM [57], POMO [33], Sym-NCO [28], BQ-NCO [16], MVMoE [60], MatNet [34], S2V-DQN [27] and COMPASS [12]. These models are all specialized and trained in single-task mode. While all these models are usually combined with a search procedure, such as beam search [16], simulation-guided beam search  [14], instance augmentation [33], active search [23], for a fair comparison of the underlying neural models, and for sake of space, we only report here the results with a greedy application of each model.
4.1 Performance on the Training Tasks
[…]
ATSP100 CVRP100 CVRPTW100 OP100
opt. gap time opt. gap time opt. gap time opt. gap time
Oracle solver 0.00% 29s 0.00% 12m 0.00% 10m 0.00% 1.1m
AM greedy - 7.11% 1s 5.35% 1s
MDAM greedy - 4.84% 1s - 2.88% 16s
POMO no aug - 1.21% 1s - -
Sym-NCO greedy - 3.33% 1s - 2.03% 2s
BQ-NCO greedy 1.27% 2s 2.79% 2s - 0.22% 2s
MVMoE/4E - 1.65% 1s 4.90% 1s -
MatNet greedy 0.93% 1s - - -
GOAL single-task greedy 0.29% 5s 2.10% 5s 2.54% 5s 0.08% 3s
GOAL multi-task greedy 0.45% 5s 2.85% 5s 3.21% 5s 0.51% 3s
KP100 MVC100 UMSP100x10 JSSP10x10
opt. gap time opt. gap time opt. gap time opt. gap time
Oracle solver 0.00% 1s 0.00% 2m 0.00% 2m 0.00% 12m
POMO no aug 0.19% - - -
BQ-NCO greedy 0.10% - - -
S2V-DQN - 0.97% 2s - -
COMPASS - - - 4.70% 23m
GOAL single-task greedy 0.10% 3s 0.31% 3s 1.82% 7s 2.65% 6m
GOAL multi-task greedy 0.12% 3s 0.34% 3s 2.28% 7s 3.23% 6m
Table 1: Performance on 128 test instances from the training tasks. Lower is better.
Generalization to new instance distributions.
Since GOAL is trained with supervised learning on instances of a certain size, it is critical that its performance generalizes to larger instances. Table 2 presents the generalization to instances up to 10 times bigger than the training ones, for a subset of the training tasks. We see an excellent generalization performance on ATSP and CVRP. For the JSSP, the generalization is not as strong but still very good on most distributions. Note that scaling to much larger instances is not a focus of this paper. Recent approaches which do not rely on quadratic attention blocks such as [39] would be more appropriate, although highly specialized to some problems. We believe that the more expensive mixed-attention blocks are crucial for the model’s capacity to handle diverse tasks. We also tested finetuning GOAL to other distribution shifts, namely to the Euclidian distance for TSP and asymmetric distances for CVRP (while GOAL was trained on ATSP and Euclidian CVRP). Results are presented in Appendix B and demonstrate a strong performance.
Train. distribution Generalization
ATSP100 ATSP200 ATSP500 ATSP1000
MatNet greedy 1.57% 1s 97.21% 12s - -
BQ-NCO greedy 1.63% 1s 1.55% 9s 4.29% 26s 8.09% 1m
GOAL multi-task greedy 0.45% 5s 1.54% 30s 1.48% 3m 1.89% 1h
CVRP 100 CVRP200 CVRP500 CVRP1000
AM greedy 7.11% 1s 10.16% 1s 17.2% 1s 72.95% 2.1s
MDAM greedy 4.84% 1s 7.54% 6s 12.94% 24s 32.25% 1m
POMO no aug 1.21% 1s 5.94% 6s 31.36% 32s 284.58% 1m
Sym-NCO greedy 3.33% 1s 8.42% 3s 37.58% 54s 481.53% 9m
BQ-NCO greedy 2.79% 1s 2.81% 9s 3.64% 55s 5.88% 2m
GOAL multi-task greedy 2.85% 5s 2.80% 12s 3.63% 48s 6.63% 3m
JSSP10x10 JSSP15x15 JSSP20x15 JSSP20x20
COMPASS 4.70% 23m 8.00% 38m 10.40% 1h -
GOAL multi-task greedy 2.65% 6m 5.90% 18m 17.40% 32m 10.80% 59m
Table 2: Generalization performance on larger instances. Lower is better.
4.2 Fine-Tuning to New Tasks
In this section, we want to evaluate the transfer capability of the trained GOAL backbone to unseen tasks. We consider four new CO problems: the Traveling Repairman Problem (TRP), the Open Vehicle Routing Problem (OVRP), the Split Delivery Vehicle Routing Problem (SDVRP) and the Maximum Independent Set (MIS) problem (detailed description in Appendix A). For each problem, our first attempt was to define new task-specific input and output adapters and train them either with or without back-propagation through the backbone. It turned out that in that case, the fine-tuning process is very unstable and has a strong dependency on the parameter initialization (i.e. the random seed). Next, we initialized the input adapters with the ones of the closest training problems: TSP for TRP, CVRP for OVRP and SDVRP and MVC for MIS. The output adapters were still trained from scratch. In this case the fine-tuning is very effective. In Fig 2, we see that the finetuned model reaches very good optimality gaps (at most $5\%$) much faster than a single-task GOAL model trained from scratch. Our few-shot fine-tuning process relies on only 64 instances and lasts less than 1 minute. We use the same loss and optimizer as for the training but with a smaller initial learning of 0.0001. These results assume that the whole model is fine-tuned, including the backbone. When the latter is frozen, we observe a slight degradation of the results with no significant speedup. From these experiments, we hypothesize that the input adapters are the difficult part of the transfer. To test this hypothesis, we considered one of the training tasks, discarded its input and/or output adapters and tried to re-learn them from scratch, with the trained backbone. This led to similar observations: re-learning the output adapter works, while re-learning the input adapter is very unstable. We conclude that transfer in GOAL is successful for problems that share the same instance structure as one of the training problems and thus can benefit from a pre-trained input adapter. Note that the new problems can have arbitrary objective and constraints (as is the case with the four tested new tasks). Appendix D further discusses input adapters and their impact on the latent representations.
Figure 2: Fine tuning GOAL to new tasks.
4.3 Ablation Study
Figure 3: Comparison of the model performance when using different variants of mixed-attention blocks for solving the Asymmetric Traveling Salesman problem
Figure 4: Comparison between single and multi-type architectures for the Job Shop Scheduling Problem
Mixed-attention blocks.
To evaluate the effectiveness of our mixed-attention blocks, we replace them by alternative modules which also aim at mixing the edge information with the attention scores, namely MatNet [34] and G2G [20]. More precisely we focus on the ATSP and train three versions of GOAL with the three different blocks. For MatNet, we keep the original hyper-parameters used for solving the ATSP. Figure 4 presents the models performance (optimality gap on a test set of 128 instances) at different points of the training. We see that using our mixed-attention blocks leads to a significant gain, on the performance at convergence (MatNet) or on the speed of convergence (G2G). Note that these ablations are performed on a smaller training dataset than in the previous experiments (100k instances), hence the wider opt gap.
Multi-type transformer.
The proposed multi-typed transformer architecture is meant to deal with tasks that are naturally represented by multipartite graphs with heterogeneous nodes or edges, typically scheduling problems where jobs and machine nodes play different roles. To evaluate its effect on the performance, we compare it to that obtained with the more direct alternative which ignores node heterogeneity: we simply concatenate the features of the different types of nodes (or edges) and use dummy values whenever a feature is undefined, so that the original single-type model can be used. We implemented this alternative for the Job Shop Scheduling Problem and present in Fig 4 a comparison between the training performance of the single-type and multi-type architectures. Clearly the multi-type architecture converges faster and to a better performance: after 200 epochs of training on 100K instances, it reaches a test optimality gap of 4.6% while the single-type model achieves only 8.9%.
5 Discussion and Conclusion
In this paper, we propose a generic neural CO model (GOAL) and demonstrate the feasibility of learning a shared backbone model to solve a variety CO problems, spanning routing, scheduling, packing and classic graph problems. We have shown the benefits of GOAL as (i) a single-task specialized model, providing state-of-the-art results on six CO problems; (ii) a multi-task model, able to solve several problems with the same backbone; and (iii) a pretrained model, to be efficiently fine-tuned for new problems. To the best of our knowledge, this is the first neural approach that handles such a variety of problems with a single backbone model. We believe this is a key step in the direction of a foundation model for CO, where the diversity of problem representations (as opposed to the uniformity of textual or visual data) is still a major hurdle.
Our experiments have exclusively been conducted with Imitation Learning, which assumes the availability of good-quality solutions for at least a certain size of instances. While this is not an intrinsic limitation of the GOAL model – the underlying MDPs being defined independently of the training process – we believe that Imitation Learning is an effective training paradigm for learning multi-task models. This is indeed demonstrated by previous works which relied on sequence-modeling training tasks for control [51, 47]. Leveraging solutions to classical NP-hard problems, provided by traditional solvers, which are based on decades of OR research, is particularly advantageous for pretraining a model. Such a model can later be adapted to more exotic problems with minimal supervision data, or even better, with Reinforcement Learning. Indeed using Reinforcement Learning to fine-tune GOAL to new problems, for which high-quality solvers may not be available, constitutes a promising direction for future work.
Appendix A Description of the considered CO problems
In this Section, we describe the CO problems that are solved by GOAL as well as the instance generation process. We organize the problems into three categories: Routing, Scheduling and Packing and Graph Problems. We also provide an overview Table of the problems and used solvers for each category.
A.1 Routing Problems
TSP TRP OP
Node features - - Price
Edge features Distance between nodes
Instance features - - Maximum tour length
Objective $min$(tour length) $min$(sum of waiting times) $max$(collected prices)
Oracle Concorde(LKH) LKH EA4OP
Action Select a node
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
CVRP/SDCVRP CVRPTW OCVRP
Node features Demands Demands Time window starts Time window ends Service times Demands
Edge features Distance between nodes
Instance features Vehicle capacity
Objective Minimize total tour length
Oracle Concorde/HGS HGS LKH
Action Select a next node, subroute is done
[Uncaptioned image] [Uncaptioned image]
Table 3: Overview of the routing problems solved by GOAL
A.1.1 Traveling Salesman Problem
The Traveling Salesman Problem (TSP) is a classic combinatorial optimization problem. The objective of the TSP is to find the shortest possible route that a salesman can take to visit a given set of cities exactly once and return to its starting point. In the path-TSP variant, the origin and destination of the salesman are distinct nodes (but may have the same coordinates, so that the TSP is a special case of the path-TSP). A TSP instance can be represented either by node coordinates or by a distance matrix. The majority of NCO models address TSP in the coordinate setting, yet the matrix setting holds greater significance for practical applications, as it applies beyond the Euclidean context, where the matrix can represent time, energy or cost, rather than distance, and need not satisfy the symmetry property of a distance matrix.
Data generation
For the Euclidian case, we generate datasets as described in [30]: points are uniformly sampled from the unit square, and the corresponding distance matrix $A$ is computed. For non-Euclidean problems (ATSP) we generate data as described in [34], where distance matrices are generated randomly, but post-processed to satisfy the triangular inequality.
A.1.2 Capacitated Vehicle Routing Problem
The Capacitated Vehicle Routing Problem (CVRP) is another standard routing problem. The goal of CVRP is to optimize the delivery routes of a fleet of vehicles to a set of customers, while considering each vehicle’s capacity. The objective is to minimize the overall distance traveled by the vehicles, while respecting the capacity constraints and ensuring that all the customer demands are met.
Data generation.
The matrix is sampled in the same way as for TSP, while node demands are randomly generated integers from the interval [1, 10]. Following [30], the vehicle capacity depends on the problem size: capacity of 50 for a problem size of 100, 80 for 200, 100 for 500, and capacity of 250 for problems with 1000 nodes.
A.1.3 CVRP with Time Windows
The CVRP with Time Windows (CVRPTW) is an extension of the CVRP. In addition to demand constraints, each customer also has a specified time window during which they can be visited, along with the necessary service time required for serving that customer. The objective of the CVRPTW is to find routes for the vehicles that satisfy the capacity limits, deliver goods within the specified time windows, and minimize the total travel and waiting time or the number of vehicles used. In our model, we consider the standard version of the problem where total traveling and serving time should be minimized.
Data generation.
For the time windows and service times, we use a similar procedure to [17], where the travel times between nodes are sampled in the range [0, $\sqrt{2}\pm 0.1$]. We ensure the feasibility of the resulting instances by sampling time windows using a more refined procedure. Windows start time values are sampled from the range (0.01, 0.25). It’s important to note that windows start time do not directly impact problem feasibility, as vehicles can wait at the node until the service starts. On the other hand, the choice of the windows end time has an impact on feasibility. An instance is unfeasible if the window end time of a customer is less than the travel time from the depot plus the service time. To avoid this, windows end time are sampled from the interval [max(window start time, travel time from depot) + 2 $\times$ service time, max(window start time, travel time from depot) + 4 $\times$ service time].
A.1.4 Split Delivery VRP
The Split Delivery VRP (SDVRP) is similar to the CVRP with the difference that a customer can be visited by several vehicles, i.e. its demand is split between multiple vehicles.
Data generation.
We use the same data as for the CVRP (Sec A.1.2).
A.1.5 Open Vehicle Routing Problem
This is a variant of CVRP where an unbounded fleet of vehicles is available at the depot, each serving a set of customers, and they do not need to return to the depot at the end of their service.
Data generation.
We use the same data as for the CVRP (Sec A.1.2).
A.1.6 Orienteering Problem
The Orienteering Problem (OP) is a combinatorial optimization problem that involves finding the optimal route to maximize the collected prizes while visiting a set of locations and satisfying a maximal distance constraint.
Data generation.
Again we follow previous work [30]: node locations are sampled in the same way as for the TSP and CVRP and prizes are randomly sampled integers from the interval [1, 10]. The maximal distance constraint is set to 4.
A.1.7 Traveling Repairman Problem
In the travelling repairman problem (TRP), as in TSP, a set of locations is given as well as the time needed to travel between any pair of locations. Each location has a task to be performed by the repairman. The objective of the problem is to find the route that minimizes the sum of waiting times for all tasks.
Data generation.
We use the same data as for the TSP (Sec A.1.1)
A.2 Packing and Graph Problems
KP MVC MIS
Node features Price and weight -
Edge features - Adjacency matrix
Instance features Knapsack capacity -
Objective $max$(collected prices) $max$(vertex coverage) $min$(independent set)
Oracle ORTools FastWVC
Action Select a node
Table 4: Overview of the packing and graph problems solved by GOAL
A.2.1 Knapsack Problem
The Knapsack Problem (KP) is a well-known combinatorial optimization problem that involves selecting a subset of items from a given set. Each item had a weight and a value, and the goal is to maximize the cumulated value of selected items, but without exceeding a given cumulated weight threshold, known as the knapsack’s (weight) capacity.
Data generation.
Following the same procedure as in [33], node weights and values are randomly sampled from the unit interval and knapsack capacity is set to 25 for all problem sizes.
A.2.2 Minimum (Weighted) Vertex Cover Problem
The goal of the Minimum Vertex Cover Problem (MVC) is to find the smallest set of vertices in a graph, such that each edge is incident to at least one of these vertices. In the weighted version of this problem (MWVC), each vertex has an associated weight, and the goal is to select vertices with the minimum cumulated sum of weights while satisfying the vertex cover condition.
Data generation.
For this problem, we follow a similar procedure as in [27]. We generate random graphs by using Erdős–Rényi model with the critical edge probability $p$ sampled from [0.05, 0.15].
A.2.3 Maximum Independent Set
In the Maximum Independent Set (MIS), we are given an undirected graph $G=(V,E)$ and the goal is to find a subset of nodes $S\subset V$ of maximal size $|S|$ such that each pair of nodes in $S$ is not connected by an edge.
Data generation.
We use the same data generation process as for the Minimum (Weighted) Vertex Cover Problem.
A.3 Scheduling Problems
UMSP JSSP
Node features Machine: Availability time Task: Execution time, Availability time Machine: Availability time
Edge features Job-machine: Execution times Task-task: precedence constraints, $ij=1$ if tasks $i$ and $j$ belong to the same job Task-machine: $ij=1$ if task $i$ must be executed on machine $j$
Objective Minimize a makespan
Oracle HiGHS ORTools
Action Select a machine for a random task Select a task
[Uncaptioned image] [Uncaptioned image]
Table 5: Overview of the scheduling problems solved by GOAL
A.3.1 Job Shop Scheduling Problem
The Job Shop Scheduling Problem (JSSP) is a classic scheduling problem. We are given $N$ jobs and $M$ machines. Each job consists of a sequence of operations that need to be processed in a specific order (implying precedence constraints). Each operation has a given processing time. Each machine can only process one operation at a time and when an operation is started it cannot be interrupted (no preemption). A solution is a schedule, defined by the start time of each operation such that the makespan, i.e. the maximum completion time of the operations, is minimized.
Data generation.
We follow the standard procedure for generating JSSP benchmark datasets: execution times are randomly sampled from $U(1,10)$.
A.3.2 Unrelated Machine Scheduling Problem
In the Unrelated Machine Scheduling problem (UMSP), the goal is to schedule $N$ jobs on $M$ machines such that the makespan is minimized. Each job $i$ has a processing time $p_{ij}$ on each machine $j$, which is arbitrary (hence the term unrelated). There are no precedence constraints and the schedules are non-preemptive. The solution to the problem is essentially a partition of the jobs on the machines. [21] shows that this problem is NP-hard.
Data generation.
For UMSP, the processing times are randomly generated from $U(1,100)$.
Appendix B Finetuning of GOAL on Various Distribution Shifts
We consider the following distribution shifts: TSP with Euclidian distances and CVRP with asymmetric distances (while GOAL was trained on ATSP and Euclidian CVRP). Fig 5 compares finetuning GOAL on these new distributions, versus training a new single-task model from scratch. We observe that finetuning converges faster and to a better optimality gap.
Figure 5: Tuning GOAL to various distribution shifts
Appendix C GOAL Hyperparameters
The GOAL backbone consists of 9 layers with an embedding size of 128, a feed-forward dimension of 512, and ReZero normalization [3]. The mixed-attention blocks have 8 heads.
Appendix D Analysis of the latent representations
To investigate the interaction between the different CO problems, we compare the latent representations of the different tasks and their evolution through the layers of the trained GOAL model. Since these are high-dimensional vectors, we use the standard t-SNE technique [53] to visualize the representations in the 2D space. More precisely, in Fig 6 we plot the t-SNE projection (with perplexity $150$) of the node embeddings after the first, fourth, and last layer of the model, for each of the training tasks (top) and with the addition of the transfer tasks (bottom). For each task, we plot 20 full instances of size 100 (110 for scheduling problems). We observe that the representations seem clustered per problem in the different layers. We also see clear overlaps between similar problems (CVRP-SDCVRP-OVRP in green shades and TSP-TRP in blue shades) in layers 1 and 4 while they become separated in the last layer. These visualizations indicate that the problem-specific input adapters seem to map the instances to distinct areas of the embedding space. This might explain the strong dependency between the input adapters and the backbone which makes learning a new input adapter for a new task often challenging.
Figure 6: Visualization of the evolution of the node latent representations through the GOAL model for the training tasks (top) and with the addition of the transfer tasks (bottom)

### v2

@@
-GOAL: A Generalist Combinatorial Optimization Agent Learner
+GOAL: A Generalist Combinatorial Optimization Agent Learning
 Abstract
@@
 In this paper, we propose GOAL (for Generalist combinatorial Optimization Agent Learning), a generalist model capable of efficiently solving multiple COPs and which can be fine-tuned to solve new COPs.
-GOAL consists of a single backbone plus light-weight problem-specific adapters, mostly for input and output processing.
+GOAL consists of a single backbone plus light-weight problem-specific adapters for input and output processing.
 The backbone is based on a new form of mixed-attention blocks which allows to handle problems defined on graphs with arbitrary combinations of node, edge and instance-level features.
-Additionally, problems which involve heterogeneous nodes or edges, such as in multi-partite graphs, are handled through a novel multi-type transformer architecture, where the attention blocks are duplicated to attend only the relevant combination of types while relying on the same shared parameters.
-We train GOAL on a set of routing, scheduling and classic graph problems and show that it is only slightly inferior to the specialized baselines while being the first multi-task model that solves a variety of COPs.
-Finally we showcase the strong transfer learning capacity of GOAL by fine-tuning or learning the adapters for new problems, with only few shots and little data.
+Additionally, problems which involve heterogeneous types of nodes or edges are handled through a novel multi-type transformer architecture, where the attention blocks are duplicated to attend the meaningful combinations of types while relying on the same shared parameters.
+We train GOAL on a set of routing, scheduling and classic graph problems and show that it is only slightly inferior to the specialized baselines while being the first multi-task model that solves a wide range of COPs.
+Finally we showcase the strong transfer learning capacity of GOAL by fine-tuning it on several new problems.
+Our code is available at https://github.com/naver/goal-co/.
 1 Introduction
-Combinatorial Optimization problems are extensively studied in the Operations Research community, because they are both challenging in terms of theoretical complexity and central in many applications such as transportation, logistics and finance.
-Traditional methods include exact solvers (typically based on integer programming [32]) and approximation algorithms [55] which provide optimality or approximation guarantees, respectively, but do not scale; and heuristic algorithms [9] which produce good-quality solutions faster but without guarantees.
+Combinatorial Optimization (CO) problems are extensively studied in the Operations Research community, because they are both challenging in terms of theoretical complexity and central in many applications such as transportation, logistics and finance.
+Traditional CO methods include exact solvers (typically based on integer programming [korte_combinatorial_2018]) and approximation algorithms [vazirani_approximation_2003] which provide optimality or approximation guarantees, respectively, but do not scale; and heuristic algorithms [boussaid_survey_2013] which produce good-quality solutions faster but without guarantees.
 Effective heuristics heavily rely on expert knowledge and are generally problem-specific.
-A more recent and promising family of methods for CO is based on machine learning, either to learn the hyper-parameters of classic CO solvers, or to directly learn heuristics in an end-to-end fashion [6, 11].
-In particular constructive neural heuristics have been successful for solving a variety of CO problems, including routing [5, 42, 30, 33], scheduling [58, 34] and graph problems [27, 45].
+A more recent and promising family of methods for CO is based on Machine Learning, either to learn the hyper-parameters of classic CO solvers, or to directly learn heuristics in an end-to-end fashion [bengio_machine_2021, cappart_combinatorial_2021].
+In particular constructive neural heuristics have been successful at solving a variety of CO problems, including routing [bello_neural_2017, nazari_reinforcement_2018, kool_attention_2019, kwon_pomo_2020], scheduling [zhang_learning_2020, kwon_matrix_2022] and graph problems [khalil_learning_2017, qiu_dimes_2022].
 However, in most cases, a separate model is specialized and trained for each problem, and sometimes even for each distribution of instances within a problem, especially for varying instance sizes.
-While much effort has been devoted to improving the generalization to different instance distributions [41, 59, 49, 16], generalization to different problems is less studied.
+While much effort has been devoted to improving the generalization to different instance distributions [manchanda_generalization_2022, zhou_towards_2023, son_meta-sage_2023, drakulic_bq-nco_2023], generalization to different problems is less studied.
 Indeed, although there are strong neural heuristics for many CO problems, any variation of these problems requires adjustment of their model and re-training from scratch.
-While it is true that some problem-specific model tailoring is always required, since the structure of problem instances and solutions are very diverse, many problems are related and it would be interesting to exploit their commonalities.
-In this paper, we seek to answer the following question: Can one model be pretrained in order to be easily specialized to solve a variety of CO problems?
-The idea of pretraining a unified model on some pretext task so that it can be efficiently fine-tuned to several downstream tasks has been extremely successful in the language and vision domains [15, 10, 46].
+While it is true that some problem-specific model tailoring may be required, since the structure of problem instances and solutions are very diverse, many problems are related and it would be interesting to exploit their commonalities.
+In this paper, we seek to answer the following question: Can we design and train a single generalist model so that it can be efficiently adapted to solve a variety of CO problems?
+The idea of pretraining a unified model on some pretext task so that it can be efficiently fine-tuned to several downstream tasks has been extremely successful in the language and vision domains [devlin_bert_2019, brown_language_2020, radford_learning_2021].
 In these domains, a self-supervised pretraining task exploits the intrinsic structure and regularities of images and text to learn generic representations.
-In contrast, in CO, while most problem instances can be modeled as graphs, there is no generic regularity to leverage by a self-supervised task.
-Graphs, unlike images or text, are a “poor” modality, with little redundancy or intrinsic semantics.
-In this paper, we propose GOAL (for Generalist combinatorial Optimization Agent Learning), a generalist model capable of efficiently solving a variety of CO problems and which can be fine-tuned to solve new related problems.
-GOAL consists of a single backbone and light-weight problem-specific adapters for input and output processing.
-To handle the large discrepancy between CO problems, we propose two complementary strategies.
-First, the backbone is based on a new form of mixed-attention blocks which allows to handle problems defined on graphs with arbitrary combinations of node, edge and instance-level features.
-Additionally, problems which involve heterogeneous nodes or edges, which typically arise with multipartite graphs, are handled through a novel multi-type transformer architecture, where the backbone mixed-attention blocks are duplicated to attend each relevant combination of types while relying on the same shared parameters.
-GOAL is able to handle any CO problem whose instances can be represented as a graph and where feasible solutions can be constructed iteratively, as a sequence of node selections.
+In contrast, in CO, while most problem instances can be modeled as graphs, there is no generic regularity to leverage by a pretext task.
+Graphs, unlike images or text contain little redundancy or intrinsic semantics.
+In this paper, we propose a new model, named GOAL (for Generalist combinatorial Optimization Agent Learning), which consists of a single backbone and light-weight problem-specific adapters for input and output processing.
+To handle the large discrepancy between CO problems, we propose three complementary strategies.
+First, to ensure learning related representations for different CO problems, we propose to use a problem-specific projection of the input instance features to a small fixed-size representation, which is then projected into the main embedding space using a shared codebook.
+Second, to handle problems defined on graphs with arbitrary combinations of node, edge and instance-level features, we design a new form of mixed-attention blocks for the backbone, injecting edge information at the core of each attention block.
+Finally, we introduce a novel multi-type transformer architecture to deal with problems which involve heterogeneous types of nodes or edges (typically represented as multipartite graphs).
+The idea is to duplicate the backbone’s mixed-attention blocks in order to attend each meaningful combination of types while relying on the same fixed set of shared parameters.
+In principle, GOAL is able to handle any CO problem whose instances can be represented as a graph and where feasible solutions can be constructed iteratively, as a sequence of node selections.
 We find that this pattern is generic enough to cover a wide range of problems.
-We train GOAL in a multi-task setting, by imitation of expert trajectories of eight classic CO problems, spanning multiple domains: Routing problems: the Asymmetric Traveling Salesman Problem (ATSP), the Capacitated Vehicle Routing Problem (CVRP), the CVRP with Time Windows (CVRPTW), the Orienteering Problem (OP); Scheduling problems: the Job Shop Scheduling Problem (JSSP) and the Unrelated Machine Scheduling Problem (UMSP); Packing problems: the Knapsack Problem (KP); and Graph problems: the Minimum Vertex Covering Problem (MVC).
+We train GOAL in a multi-task setting, by imitation of expert trajectories of eight classic CO problems, spanning multiple domains: Routing problems: the Asymmetric Traveling Salesman Problem, the Capacitated Vehicle Routing Problem with or without Time Windows, the Orienteering Problem ; Scheduling problems: the Job Shop Scheduling Problem and the Unrelated Machine Scheduling Problem; Packing problems: the Knapsack Problem; and Graph problems: the Minimum Vertex Covering Problem.
 The expert trajectories are based on good-quality solutions, easily generated by traditional problem-specific solvers for instances of medium size (generally 100 nodes).
 Experimentally we first show that GOAL performs very well on all the training tasks, making it the first model that solves such a variety of CO problems.
-Second, we illustrate GOAL’s strong zero (or few) shot generalization performance to various distributions shifts from the training tasks’ instances.
-Third, we fine-tune GOAL on four new problems and demonstrate that it achieves a good performance quickly using little data.
-We also discuss scenarios where fine-tuning is less effective, hinting at the required conditions for successful transfer; namely that the new task is close enough to at least one of the training tasks to share a similar instance structure, while its objective and constraints can be completely different.
-We further provide an ablation study to evaluate the effectiveness of our multi-type transformer architecture and mixed-attention blocks.
-Notably our architecture gives state-of-the-art results on 6 of our 8 training tasks when used in a single-task version of GOAL.
+Second, we consider eight new problems and demonstrate that GOAL can be fine-tuned to reach a good performance, either in a few minutes using few-shot imitation learning or in a couple of hours without supervision.
+In both cases, the fine-tuning takes a fraction of the time and data needed to train a model from scratch.
+We further provide an ablation study to evaluate the effectiveness of a shared codebook, multi-type transformer architecture and mixed-attention blocks.
+Notably the GOAL architecture gives state-of-the-art results on 7 out of our 8 training tasks when used in as single-task model.
 In summary, our contributions are the following:
   • We introduce GOAL: a generic and flexible model for end-to-end combinatorial optimization.
-Two key ingredients of GOAL are the simple yet effective form of mixed-attention blocks which enables handling problems represented by graphs with any combination of node, edge and instance-level features and its multi-type transformer architecture which allows to seamlessly tackle problems with multipartite graph structures.
+Two key ingredients of GOAL are the simple yet effective form of mixed-attention blocks which enables handling problems represented by graphs with any combination of node-, edge- and instance-level features and its multi-type transformer architecture which allows to seamlessly tackle problems with multipartite graph structures.
   • We demonstrate that it is possible to learn a single backbone model to effectively solve a variety of CO tasks, spanning routing, scheduling, packing and graph problems.
-  • We showcase the potential of such a multi-task pretrained model to be efficiently adapted to handle new CO tasks and provide empirical insights on how closely related the new tasks should be to the training tasks to ensure successful transfer.
+  • We showcase the potential of such a multi-task pretrained model to be efficiently adapted to handle varied new CO tasks, with or without supervision.
+We believe this is a key step toward a foundation model for CO.
 2 Related Works
@@
 Our work falls into the category of end-to-end constructive heuristics for combinatorial optimization.
-Constructive neural heuristics have been successfully applied to a wide range of COPs, including many variants of routing problems [30, 57, 28, 31, 18, 52], scheduling problems [58, 34], as well as classic COPs such as the Knapsack Problem [33] and the Maximum Independent Set [27, 52].
-Another successful category of neural CO methods are improvement heuristics, which focus on iteratively enhancing a given complete solution, typically by using a neural network to guide the selection of local search operators [13, 22, 40, 56].
-The seminal work of [27] introduces graph embeddings which do work for a variety of CO problems but with limited performance.
+Constructive neural heuristics have been successfully applied to a wide range of COPs, including many variants of routing problems [kool_attention_2019, xin_multi-decoder_2021, kim_sym-nco_2022, kool_deep_2021, fu_generalize_2021, sun_difusco_2023], scheduling problems [zhang_learning_2020, kwon_matrix_2022], as well as classic COPs such as the Knapsack Problem [kwon_pomo_2020] and the Maximum Independent Set [khalil_learning_2017, sun_difusco_2023].
+Another successful category of neural CO methods are improvement heuristics, which focus on iteratively enhancing a given complete solution, typically by using a neural network to guide the selection of local search operators [chen_learning_2019, hottung_neural_2019, ma_learning_2021, wu_learning_2020].
+The seminal work of [khalil_learning_2017] introduces graph embeddings which do work for a variety of CO problems but with limited performance.
 Current state-of-the-art models are tailored for certain problem structures (e.g.
-Euclidian routing problems, where nodes are represented by their coordinates plus some extra node or graph features such as demand or vehicle capacity), and successfully handled by transformer-based architectures [30].
-For non-euclidian routing problems, [34] and [16] propose adapted architectures to deal with the distance matrix, but they are specialized for such problems and do not apply to those with node-only features.
-In GOAL, we propose a novel mixed-attention block which can flexibly deal with both node and edge features, incurring no overhead when one or the other is missing.
+euclidian routing problems, where nodes are represented by their coordinates plus some extra node or graph features such as demand or vehicle capacity), and successfully handled by transformer-based architectures [kool_attention_2019].
+For non-euclidian routing problems, [kwon_matrix_2022] and [drakulic_bq-nco_2023] propose adapted architectures to deal with the distance matrix, but they are specialized for such problems and do not apply to those with node-only features.
+In GOAL, we propose a novel mixed-attention modules which can flexibly deal with both node and edge features, incurring no overhead when one or the other is missing.
 Generalization of neural heuristics
-While the first seminal models for neural CO had a dramatic drop in performance when tested on out-of-training distribution instances [26, 41], several recent approaches show a greatly improved generalization performance, especially to larger instances [18, 35, 49, 43], to distribution shifts of other instance parameters [8], or both [41, 8, 59, 16].
+While the first seminal models for neural CO had a dramatic drop in performance when tested on out-of-training distribution instances [joshi_learning_2022, manchanda_generalization_2022], several recent approaches show a greatly improved generalization performance, especially to larger instances [fu_generalize_2021, li_learning_2021, son_meta-sage_2023, pan_h-tsp_2023], to distribution shifts of other instance parameters [bi_learning_2022], or both [manchanda_generalization_2022, bi_learning_2022, zhou_towards_2023, drakulic_bq-nco_2023].
 Concurrently with our work, some papers have started to explore the cross-problem generalization for the family of routing problems.
-Specifically, [36] propose to pretrain the AM model [30] on the TSP and to fine-tune either the entire model or some light-weight modules on the Price Collecting TSP and OP.
-Additionally, [60] and [37] explore generalization to different variants of the CVRP.
-They build on the idea of zero-shot compositionality [48] by training a model on CVRPs with some combinations of constraints from a predefined set (e.g.
+Specifically, [lin_cross-problem_2024] propose to pretrain the AM model [kool_attention_2019] on the TSP and to fine-tune either the entire model or some light-weight modules on the Price Collecting TSP and OP.
+Additionally, [zhou_mvmoe_2024, liu_multi-task_2024, berto_routefinder_2024] explore generalization to different variants of the CVRP.
+They build on the idea of zero-shot compositionality [ruis_independent_2021] by training a model on CVRPs with some combinations of constraints from a predefined set (e.g.
 time windows and duration limit constraints) and show that the model generalizes to some unseen combinations of the same constraints.
-Although these approaches tackle multiple problems, it is important to note that the structures of the considered Euclidian routing problems are very similar and indeed many of those problems were solved by the seminal Attention Model of [30] with little adaptations of the base model.
-Our paper goes further and considers multi-task training and transfer learning of a new model on a variety of COPs with different structure.
+Although these approaches tackle multiple problems, it is important to note that the considered Euclidian routing problems remain very similar in structure and indeed many of those problems were solved by the seminal Attention Model of [kool_attention_2019] with little adaptations of its base model.
+Our paper goes further and considers multi-task training and transfer learning of a new model on a variety of COPs with different structures.
 Multi-task training for sequential decision making
 Multi-task (pre-)training has recently shown promising results in related fields.
-For example, [25] proposed a graph neural network processor capable of learning to execute a variety of algorithms, such as sorting and searching.
+For example, [ibarz_generalist_2022] proposed a graph neural network processor capable of learning to execute a variety of algorithms, such as sorting and searching.
 The paper focuses only on polynomial problems and learns to imitate the execution of classical algorithms.
-In contrast, we address NP-hard problems and learn by imitation of the solution provided by (potentially exponential) oracle algorithms.
-Recent approaches in control rely on sequence modeling of expert trajectories from multiple tasks [50, 51].
-Although they treat various tasks, with potentially diverse state and action spaces, these works still consider image observations and exploit this structure.
-[47] proposes GATO, a general-purpose agent obtained by training a large transformer-based model on a hundreds of tasks spanning vision, language and control.
+In contrast, we address NP-hard problems and learn by imitation of the solution provided by (potentially exponential) specialized solvers.
+Recent approaches in control rely on sequence modeling of expert trajectories from multiple tasks [stooke_decoupling_2021, sun_smart_2023].
+Although they treat various tasks, with potentially diverse state and action spaces, these works still consider image observations and exploit this common structure.
+[reed_generalist_2022] proposes GATO, a general-purpose agent obtained by training a large transformer-based model on hundreds of tasks spanning vision, language and control.
 It demonstrates the potential of a single network (albeit with 1.2B parameters) to effectively tackle a wide variety of tasks.
@@
 Figure 1: Left: Architecture of GOAL for single-type problems; the green components are task-specific, while the others belong to the backbone.
-Right: Architecture of GOAL for multi-type problems (with here 2 types of nodes T1,T2); “self” and “cross” are identical multi-head mixed attention blocks (same parameters for all types and tasks); “FF” for the two types are also identical Feed Forward blocks; for the edges, all the type combinations are shown (here T11,T12,T21,T22) but each task can define its own.
+The codebook is a shared $L{\times}D$ (resp.
+$\bar{L}{\times}\bar{D}$) matrix applied to the node (resp.
+edge) representations produced by the input adapters.
+Right: Architecture of GOAL for multi-type problems (with here 2 node types T1,T2); “self” and “cross” are identical multi-head mixed attention blocks (same parameters per layer for all types and tasks); “FF” for the two types are also identical Feed Forward blocks per layer.
+For the edges, all the type combinations are shown (here T11,T12,T21,T22) but each task can define its own.
 The same set of parameters is used for the two architectures.
@@
 an element of $\arg\min_{x\in X}f(x)$.
-Constructive approaches assume that $f$ is defined not only over $X$ but over a set of partial solutions, which contains $X$ and is amenable to a step-wise construction procedure: each partial solution can be obtained by a sequence of construction steps.
+Constructive approaches assume that $f$ is defined not only over $X$ but over a set of partial solutions, which includes $X$ and is amenable to a step-wise construction procedure where each partial solution can be obtained by a sequence of construction steps.
 Thus, solving the problem instance can be formulated as a sequential decision procedure which iteratively selects the construction steps so as to reach not only a feasible but an optimal solution.
 That in turn can be modeled as a Markov Decision Problem (MDP) associated with the problem instance.
-The action space of this MDP is naturally defined as the (discrete) set of construction steps, but there are many ways to choose its state space.
-For the very broad class of so called “tail-recursive” CO problems, [16] shows that an effective choice is the space of all possible instances of the problem.
-Indeed, the construction of a solution can be modeled as a trajectory in the space of problem instances, following the tail sub-problems [7] obtained after each construction step.
-All the classical standard NP-hard CO problems, such as vehicle routing problems, machine scheduling, various graph problems etc.
-are tail-recursive, sometimes modulo a slight reformulation (e.g.
-TSP is not per se tail-recursive, but path-TSP is).
-In the sequel, we deal exclusively with tail-recursive CO problems and their construction MDPs, called BQ-MDP in [16].
+In this paper, we use the BQ-MDPs introduced by [drakulic_bq-nco_2023] which are defined for the widespread class of tail-recursive problems  [bertsekas_dynamic_vol1_2012].
+In BQ-MDPs, the actions are the construction steps, typically the selection of a node, while the states are instances.
 3.2 Multi-Task CO Problems
 Now, given a class $\mathcal{T}$ of tail-recursive CO problems called tasks, let $\boldsymbol{\Omega}_{t}$ for each task $t{\in}\mathcal{T}$ denote the parametric space used to represent its instances.
-Consider the new (multi-task) CO problem defined as the union of the tasks in $\mathcal{T}$ (each being a set of instances).
-Its instances can be represented in the disjoint union $\biguplus_{t\in\mathcal{T}}\boldsymbol{\Omega}_{t}$ consisting of pairs $(t,\omega)$ of a task identifier $t{\in}\mathcal{T}$ and a point $\omega{\in}\boldsymbol{\Omega}_{t}$ in its corresponding parameter space.
+Consider the new (multi-task) CO problem whose instances are represented in the disjoint union $\biguplus_{t\in\mathcal{T}}\boldsymbol{\Omega}_{t}$ consisting of pairs $(t,\omega)$ of a task identifier $t{\in}\mathcal{T}$ and a point $\omega{\in}\boldsymbol{\Omega}_{t}$ in its corresponding parameter space.
 It is easy to show that the multi-task problem is tail-recursive whenever each of its components is.
 And its BQ-MDP is then simply the disjoint union of the BQ-MDPs of its components.
-In this paper, we study alternatives for the multi-task policy space, beyond the trivial one where no parameter is shared between the policies of the different tasks (Cartesian product of the policy spaces of the individual tasks).
+In this paper, we study alternatives for the multi-task policy space, beyond the (trivial) Cartesian product of the policy spaces of the individual tasks, where no parameter is shared between the different tasks.
 The underlying assumption is that if the tasks are sufficiently similar, learning one task may elicit implicit skills useful to other tasks, which are then leveraged by parameter sharing.
 In the extreme, some tasks may require little training data and/or just a few shot fine-tuning to obtain acceptable performance, leveraging their similarity to more thoroughly trained tasks.
-Since each task has its own parametric space used to represent its instances, some part of any multi-task policy has to be task specific, and trained by processing at least a few instances of each task: true zero-shot generalization is not possible.
-The situation is close to multi-lingual Natural Language Processing, where the vocabulary of each language needs to be seen at least once at training.
-And we borrow the solution used in that context, namely adapters [4], to capture the task-specific parts of the policy.
+Since each task has its own parametric space used to represent its instances, some part of any multi-task policy has to be task specific, and trained by processing at least a few instances of each task: pure zero-shot generalization is not possible.
+The situation is close to multi-lingual Natural Language Processing, where the vocabulary of each language needs to be seen at least once at training or fine-tuning.
+And we borrow the solution used in that context, namely adapters [bapna_simple_2019].
 3.3 Task Representation
 To leverage operational similarity between tasks, we need them to have at least representational similarity.
-Thus, we assume that each instance of size $N$ of a task (CO problem) is defined by a set of $N$ abstract “nodes”, together with features which can be attached at node, edge (node pair) or instance level.
+Thus, we assume that each instance of size $N$ of a task (CO problem) is represented by a set of $N$ abstract “nodes”, together with features which can be attached at node, edge (node pair) or instance level.
 Instance features can be replicated and appended to the features of each node, hence can be ignored without loss of generality.
@@
 Wherever appropriate, task-dependent masking can be used to ensure that probabilities are null for the actions disallowed by a task.
-For example, for the path-TSP task, the nodes are the customers to visit plus the origin and destination of the vehicle.
-There are two proper node features, each holding a binary value indicating whether the node is the origin or the destination, respectively (thus $F_{t}{=}3$ counting node identifier).
-And there is a single edge feature ($\bar{F}_{t}{=}1$) holding the edge cost (distance, time, energy, etc.).
-The action is a simple node selection ($K_{t}{=}1$), with origin and destination nodes being masked.
-In a BQ-MDP transition, the selected node becomes the new origin, and the previous origin is discarded ($N$ is decreased by $1$).
 3.4 Architecture of the Model
-The architecture of our model generalizes that of BQ-NCO [16], and is described in Fig.
+The architecture of our model generalizes that of BQ-NCO [drakulic_bq-nco_2023], and is described in Fig.
 1 left.
-It essentially consists of a sequence of layers whose parameters are shared by all the tasks (backbone) together with task-specific adapter modules at the input and output of the model [^1] Additional adapters inside layers can also be introduced, but this paper considers only input-output adapters..
+It essentially consists of a sequence of layers whose parameters are shared by all the tasks (backbone) together with task-specific adapter modules at the input and output of the model [^1] Adapters can also be introduced within layers, but this paper considers only input-output adapters..
 The input-output adapter modules for task $t$ are specific to the dimensions $F_{t},\bar{F}_{t},K_{t}$ and the meaning of the corresponding features.
-No positional encoding is used (all relational information is assumed to be captured by $\bar{x}$).
+No positional encoding is used: all relational information is assumed to be captured by $\bar{x}$.
 The layers themselves, on the other hand, are only aware of the shared embedding dimensions $D,\bar{D}$.
-They are self-attention transformer layers [54] where vanilla attention is replaced by multi-head mixed attention, similar in intent to [34] but with a different design.
-Our mixed attention block, in the general case where $M$ nodes attend on $N$ nodes ($M{=}N$ in self-attention), is an instance of generalized attention [2].
-It takes the following inputs: node representations as vectors for $N$ queries, $M$ keys, and $M$ values, with an optional $M{\times}N$ log-binary mask $\mathcal{M}$, as in vanilla attention, plus optional edge representations as $M{\times}N$ “edge” vectors.
-In each head $h$, each query, key, value vector is first projected, as in vanilla attention, into a common space of dimension $d$, yielding for each node $n{=}1{:}N$ and $m{=}1{:}M$, vectors:
-$Q_{n}^{(h)}=Q_{n}\mathbf{W}_{Q}^{(h)},\hskip 14.22636ptK_{m}^{(h)}=K_{m}\mathbf{W}_{K}^{(h)},\hskip 14.22636ptV_{m}^{(h)}=V_{m}\mathbf{W}_{V}^{(h)}.$
-The edge vectors are also projected into the same space of dimension $d$, but twice so: once as a key and once as a query, resulting in vectors:
-$Q^{\prime(h)}_{mn}=E_{mn}\mathbf{W}^{\prime(h)}_{Q},\hskip 14.22636ptK^{\prime(h)}_{mn}=E_{mn}\mathbf{W}^{\prime(h)}_{K}.$
-Vanilla attention then builds an $M{\times}N$ score matrix $S_{mn}^{(h)}{=}Q_{n}^{(h)}K_{m}^{(h)\top}$, from which the attention coefficients are computed (using softmax and involving the mask).
-We simply augment the factors of the scalar product defining the score matrix with the edge projections, and keep the rest unchanged, yielding new representations $r$ of the $N$ nodes:
-$\begin{split}r=\sum_{h}\textrm{softmax}_{\textrm{col}}(S^{(h)}{+}\mathcal{M})^{\top}V^{(h)}\mathbf{W}_{O}^{(h)\top}\hskip 2.84544pt\\ \text{ with }\hskip 2.84544ptS^{(h)}_{mn}{=}(Q_{n}^{(h)}{+}Q^{\prime(h)}_{mn})(K_{m}^{(h)}{+}K^{\prime(h)}_{mn})^{\top}.\end{split}$ (1)
-In GOAL, all the node inputs are of dimension $D$, the edge inputs of dimension $\bar{D}$, and they are all projected into a space of dimension $d$ equal to $D$ divided by the number of heads.
-Furthermore, no masking is used in the attention blocks.
-One well-known limitation of this model is the quadratic complexity bottleneck of attention matrices, which are of size $N^{2}$.
-In practice though, for many tasks, a simple, task-specific, heuristic can deal with that problem.
-Essentially, instead of presenting all the $N$ nodes of the input instance to the model, the heuristic selects a priori a subset of the $n$ most promising ones.
-Thus, the attention matrices never exceed size $n^{2}$, even when $N^{2}$ would be unmanageable.
-The performance impact is often negligible, and even sometimes improving (maybe due to the dilution of attention coefficient when $N$ is large).
-For example, with TSP, only the origin, destination and 100-nearest neighbors to the origin are presented.
+They are self-attention transformer layers [vaswani_attention_2017] where vanilla attention is replaced by multi-head mixed attention, similar in intent to [kwon_matrix_2022, henderson_transformers_2023] but with a different design.
+Our mixed attention block, in the general case where $M$ nodes attend on $N$ nodes ($M{=}N$ only in self-attention), is an instance of generalized attention [andreoli_convolution_2019].
+A generalized attention mechanism takes the following inputs: node representations as vectors for $N$ queries, $M$ keys, and $M$ values, with an optional $M{\times}N$ log-binary mask $\mathcal{M}$.
+In each head $h$, each query, key, value vector is first projected into a common space of dimension $d$, yielding for each $n{=}1{:}N$ and $m{=}1{:}M$, vectors (of dimension $d$):
+$Q_{n}^{(h)}$ $=Q_{n}\mathbf{W}_{Q}^{(h)},$ $K_{m}^{(h)}$ $=K_{m}\mathbf{W}_{K}^{(h)},$ $V_{m}^{(h)}$ $=V_{m}\mathbf{W}_{V}^{(h)}.$
+Then the key and query vectors are used to compute an $M{\times}N$ score matrix $S$, which is transformed into an attention matrix (by adding the mask and applying softmax) multiplied by the reprojected value vectors to produce the output representation $r$ of the $N$ input query nodes as an $N{\times}D$ matrix:
+$r$ $=\sum_{h}\underbrace{\textrm{softmax}_{\textrm{col}}(S^{(h)}{+}\mathcal{M})^{\top}}_{\textrm{attention matrix}}V^{(h)}\mathbf{W}_{O}^{(h)\top}.$
+Attention mechanisms then differ in the way they compute the score matrix:
+[…]
+[…]
+In vanilla attention, the scores are obtained by simple scalar product of the key and query vectors ((1a)).
+Our mixed attention mechanism takes as additional input $M{\times}N$ “edge” vectors, projected twice into the common space of dimension $d$, once as key and once as query, resulting in vectors (of dimension $d$):
+$Q^{\prime(h)}_{mn}$ $=E_{mn}\mathbf{W}^{\prime(h)}_{Q},$ $K^{\prime(h)}_{mn}$ $=E_{mn}\mathbf{W}^{\prime(h)}_{K}.$ (2)
+The score matrix is still obtained as a scalar product of a key and a query vector, but now, each of them is the sum of a node component and an edge component ((1b)).
+Our mixed attention allows to incorporate arbitrary relational information (edge features) into the core of the attention mechanism.
+It differs from positional encoding, which attempts to distill a priori the edge information (limited to sequential ordering) into the node representations.
+In the GOAL backbone, all the node (resp.
+edge) inputs to the attention blocks are of dimension $D$ (resp.
+$\bar{D}$), and they are all projected into a common space of dimension $d$ equal to $D$ divided by the number of heads.
+No masking is used in the backbone.
+One well-known limitation of this model is its quadratic complexity, the attention matrices being of size $N^{2}$.
+In practice though, for many tasks, a simple, task-specific heuristic can deal with that problem by pre-filtering the set of nodes to present to the model (usually much fewer than $N$).
+Whenever available, we use such heuristics.
+Input and output adapters are linear projections, of the task-specific instance features to the backbone’s input embedding (dim.
+$D$) and from the backbone’s output embedding to task-specific action space.
+However, when $D$ is large and the number of training tasks is small, there is a risk that the embeddings produced by each task end up occupying complementary subspaces of the embedding space throughout the layers, which would defeat the purpose of effectively sharing parameters across tasks.
+To prevent this, we simply force the input linear projection to be low rank.
+Concretely, node (resp.
+edge) features are first projected into a “small” dimension space, forcing the representations to share dimensions across tasks, then this low dimension representation is plunged into the “large” $D$-dimensional embedding space using a shared codebook (see Fig.
+1 left).
 3.5 Multi-Type Transformer Architecture
 Many tasks (CO problems) involve different types of nodes.
-For example, scheduling problems often involve “operation” nodes (the things to schedule) and “machine” nodes (which constrain feasible schedules).
+For example, scheduling problems often involve “operation” and “machine” nodes.
 For GOAL, one solution is to simply treat them all indiscriminately in the backbone.
 The adapters, being problem specific, are of course free to treat them differently according to their types.
-This is sometimes unsatisfactory though, since it forces the adapter to provide uniform embeddings for all the edges, though they are often heterogeneous in the input: in the scheduling example, edge features may be precedence constraints between operations, or processing times between machines and operations, and no information at all between machines.
-Instead of packing all these different types of relations into a single big tensor, with padding values wherever a feature is not defined for a given edge, we propose a multi-type architecture where the adapters output separate embeddings per type (for nodes) or compatible pairs of types (for edges).
-Each layer of the model, instead of being a monolithic self-attention block operating on all the nodes, becomes a task-specific combination of both cross- and self-attention blocks, as described by (1), operating on different type-compatible subsets of nodes and edges.
+This is sometimes unsatisfactory though, since it forces the adapter to provide uniform embeddings for all the nodes and edges, though they are often heterogeneous by nature: in the scheduling example, edge features may carry precedence constraints when between operations, or processing times when between machines and operations, and no information at all when between machines.
+Instead of packing all these different types of relations into a single vector, with padding values wherever a feature is undefined for a given edge, we propose a multi-type architecture where the adapters output separate embeddings per type (for nodes) or compatible pairs of types (for edges).
+Each layer of the model, instead of being a monolithic self-attention block operating on all the nodes, becomes a task-specific combination of both cross- and self-attention blocks, as described by (1b), operating on different type-compatible subsets of nodes and edges.
 This is illustrated in Fig.
@@
 Therefore, in our scheduling example, the model will contain separate self-attention blocks, for operation nodes and machine nodes respectively, and cross-attention blocks, for operation-machine edges and vice-versa.
-Note that even when edges are not oriented, it is useful to have two cross-attention blocks with transposed matrices as edge embeddings: the projections in (1) are then used to interpret differently the two directions of the edges in each attention block.
-Finally, observe that while the architecture configuration is now task-specific, the model parameters remain the same for all the possible configurations.
-This is precisely what enables us to train a unique model to solve various single and multi-type problems.
+Observe that while the architecture configuration is now task-specific, the model parameters remain the same for all the configurations: all the attention blocks of a layer, be them self- or cross- attentions, and whatever the types they involve, share the same, layer-specific parameters.
+This is precisely what allows us to train a unique model to solve arbitrary single- and multi-type problems.
 4 Experiments
-Our experiments seek to answer the following questions: Does the proposed GOAL approach enable learning a single backbone to effectively solve a variety of CO problems?
-(Sec 4.1); Can we use it as a pretrained model and efficiently fine-tune it to solve new unseen CO problems?
-(Sec 4.2); and finally, to what extent do the introduced mixed-attention blocks and multi-type architecture contribute to the results?
-(Sec 4.3).
-We start by describing the multi-task training of GOAL.
+Our experiments aim to address the following questions: (i) does the GOAL approach enable learning a single backbone to effectively solve a variety of CO problems (Sec 4.1)?
+(ii) can it be used as a pretrained model and efficiently fine-tuned to solve new CO problems, unseen at training (Sec 4.2)?
+(iii) to what extent do our mixed-attention blocks and multi-type architecture contribute to the results (Sec 4.3)?
+We begin by describing the multi-task training of GOAL.
 Training tasks.
@@
 For each problem, we generate a training dataset of 1 million random instances and solve them using traditional solvers.
-Specifically, we consider the following problems and oracle solvers: the Asymmetric Traveling Salesman Problem (ATSP) with LKH [19], the Capacitated Vehicle Routing Problem (CVRP) with LKH [19], the Capacitated Vehicle Routing Problem with Time Windows (CVRPTW) with HGS [19] the Orienteering Problem (OP) with A4OP [29], the Job Shop Scheduling Problem (JSSP) with ORTools [44], the Unrelated Machine Scheduling Problem (UMSP) with HiGHS [24], the Knapsack Problem (KP) with ORTools [44] and the Minimum Vertex Covering Problem (MVC) with FastWVC [1].
-The training instances are of size 100 nodes, or 10$\times$10 for JSSP and 10$\times$100 for UMSP (resulting in both cases in 100 operation nodes and 10 machine nodes).
-For the sake of space, the description of the problems and associated data generation processes is provided in Appendix A.
+Specifically, we consider the following problems and oracle solvers: the Asymmetric Traveling Salesman Problem (ATSP) with LKH [helsgaun_extension_2017], the Capacitated Vehicle Routing Problem (CVRP) with LKH [helsgaun_extension_2017], the Capacitated Vehicle Routing Problem with Time Windows (CVRPTW) with HGS [vidal_hybrid_2022] the Orienteering Problem (OP) with A4OP [kobeaga_efficient_2018], the Knapsack Problem (KP) with ORTools [perron_or-tools_2022], the Minimum Vertex Covering Problem (MVC) with FastWVC [cai_fastwvc_2023], the Unrelated Machine Scheduling Problem (UMSP) with HiGHS [huangfu_parallelizing_2018] and the Job Shop Scheduling Problem (JSSP) with ORTools [perron_or-tools_2022].
+The training instances are of size 100 nodes, or 10$\times$10 for JSSP and 100$\times$20 for UMSP (resulting in both cases in 100 operation nodes).
+The description of the problems and the associated data generation processes is provided in Appendix E.
 Training process.
-GOAL is trained by alternating batches from each of the training tasks (round robin), using a single-class cross entropy loss for ATSP, CVRP, CVRPTW, OP, JSSP and UMSP and a multi-class cross entropy loss for KP and MVC.
-Similarly to [16], we create a batch by sampling suffix subproblems for each instance.
-We use AdamW optimizer [38], with an initial learning rate of 0.0005 and a decay rate every 10th epochs.
-We trained the model on 8 Nvidia A100 GPU servers, with batch size of 128 for 7 days, allowing the completion of approximately 500 epochs.
-The best model was selected based on the average performance on a validation set of 512 instances per problem.
-The model hyperparameters are described in Appendix C.
+GOAL is trained by alternating batches in random order from each of the training tasks, using a single-class cross-entropy loss for ATSP, CVRP, CVRPTW, OP, JSSP and UMSP and a multi-class cross-entropy loss for KP and MVC.
+Similarly to [drakulic_bq-nco_2023], we create a batch by sampling suffix subproblems for each instance.
+We use AdamW optimizer [loshchilov_decoupled_2019], with an initial learning rate of 0.0005 and a decay rate of 0.97 every 10th epoch.
+We trained the model on 8 Nvidia V100 GPU servers, with batch size of 256 for 7 days, completing ca.
+400 epochs.
+The best model was selected based on the average performance on a validation set of 128 instances per problem.
+The model hyperparameters are described in Appendix A.
 Baselines.
-We consider the following constructive neural models as baselines: AM [30], MDAM [57], POMO [33], Sym-NCO [28], BQ-NCO [16], MVMoE [60], MatNet [34], S2V-DQN [27] and COMPASS [12].
-These models are all specialized and trained in single-task mode.
-While all these models are usually combined with a search procedure, such as beam search [16], simulation-guided beam search  [14], instance augmentation [33], active search [23], for a fair comparison of the underlying neural models, and for sake of space, we only report here the results with a greedy application of each model.
+We consider the following constructive neural models as baselines: AM [kool_attention_2019], MDAM [xin_multi-decoder_2021], POMO [kwon_pomo_2020], Sym-NCO [kim_sym-nco_2022], BQ-NCO [drakulic_bq-nco_2023], MVMoE [zhou_mvmoe_2024], MatNet [kwon_matrix_2022], S2V-DQN [khalil_learning_2017] and COMPASS [chalumeau_combinatorial_2023].
+All models except MVMoE are specialized and trained in single-task mode.
+They are often combined at test time with various enhancers, such as beam search [drakulic_bq-nco_2023], simulation-guided beam search  [choo_simulation-guided_2022], instance augmentation [kwon_pomo_2020], or active search [hottung_efficient_2022], but for fair comparison, all the results we report are obtained by a pure greedy application of the trained policy at test time.
 4.1 Performance on the Training Tasks
-[…]
+In Table 1 we compare the performance of GOAL to the relevant neural baselines for each problem as well as its single-task version (i.e.
+GOAL trained separately for each task).
+First, we observe that the best performance is achieved by the single-task GOAL in 7 out of the 8 tasks, with a computation time comparable to that of the baselines.
+This proves the versatility and effectiveness of our proposed architecture, even for a specialized model.
+Second, we see that multi-task GOAL is only slightly worse than the single-task version.
+These results validate the capability of the GOAL backbone, together with the task-specific adapters, to effectively solve the variety of training tasks.
+More broadly, they demonstrate that multi-task learning can be a successful paradigm for neural combinatorial optimization.
+In addition, in Appendix D, we provide experimental results demonstrating GOAL’s excellent zero-shot generalization on instances up to 10 times larger than the training instances.
+We can note in particular the strong generalization for ATSP which seems much more challenging for the state-of-the-art baselines.
+This confirms the effectiveness of our mixed-attention mechanism.
 ATSP100 CVRP100 CVRPTW100 OP100
@@
 MatNet greedy 0.93% 1s - - -
-GOAL single-task greedy 0.29% 5s 2.10% 5s 2.54% 5s 0.08% 3s
-GOAL multi-task greedy 0.45% 5s 2.85% 5s 3.21% 5s 0.51% 3s
-KP100 MVC100 UMSP100x10 JSSP10x10
+GOAL single-task greedy 0.21% 10s 2.31% 10s 2.79% 10s -0.35% 3s
+GOAL multi-task greedy 0.50% 10s 3.06% 10s 2.86% 11s 0.28% 3s
+KP100 MVC100 UMSP100x20 JSSP10x10
 opt. gap time opt. gap time opt. gap time opt. gap time
@@
 COMPASS - - - 4.70% 23m
-GOAL single-task greedy 0.10% 3s 0.31% 3s 1.82% 7s 2.65% 6m
-GOAL multi-task greedy 0.12% 3s 0.34% 3s 2.28% 7s 3.23% 6m
+GOAL single-task greedy 0.10% 3s 0.26% 3s 1.30% 7s 2.70% 15s
+GOAL multi-task greedy 0.14% 2s 0.37% 9s 2.01% 7s 3.74% 15s
 Table 1: Performance on 128 test instances from the training tasks. Lower is better.
-Generalization to new instance distributions.
-Since GOAL is trained with supervised learning on instances of a certain size, it is critical that its performance generalizes to larger instances.
+4.2 Fine-Tuning to New Tasks
+In this section, we want to evaluate the transfer capability of the trained GOAL backbone to unseen tasks.
+We consider eight diverse new COPs: the Traveling Repairman Problem (TRP), the Prize Collecting TSP (PCTSP), the Open CVRP (OCVRP), the Split Delivery CVRP (SDCVRP), the Sequential Ordering Problem (SOP), the Maximum Covering Location Problem (MCLP), the Maximal Independent Set (MIS) problem and the Open Shop Scheduling Problem (OSSP) (detailed descriptions in Appendix E).
+GOAL can be fine-tuned to specific problems in two modes.
+In both, the adapters are trained from scratch, and the backbone is also open to backpropagation (no parameter freeze).
+The supervised mode is identical to training, except it processes only a few (labeled) instances of the problem for a few steps.
+Details and results are given in Appendix B.
+The unsupervised mode is more challenging, but also more realistic, since we may not have access to good-quality solutions for the new problem.
+We propose to fine-tune the model by iterative imitation-improvement steps between an Apprentice and an Expert, as in ExIt [anthony_thinking_2017].
+While Exit relies on a possibly complex tree-search for improvement, we use instead a brute-force but fast method, where the Expert samples multiple solutions (in our case 128) per instance from the Apprentice and selects the best, if it is sufficiently better, and sends it back to the Apprentice (current model) for imitation.
+This is similar to the self-labeling strategy proposed by corsini_self-labeling_2024 for the JSSP.
+One could also use more sophisticated experts such as generic local search heuristics or smarter sampling pirnay_self-improvement_2024 to accelerate the finetuning.
+Performance results are given in Fig 3, which shows that unsupervised fine-tuning systematically outperforms unsupervised training of the whole model from scratch at a comparable processing time budget.
+We also compare finetuning GOAL to finetuning the single-task models and report the results in Fig 2.
+We observe that the multi-task model consistently performs as well or better than the “most-related” single-task model (e.g.
+CVRP model for OCVRP or JSSP model for OSSP), hinting to the synergistic effect of the multi-task training.
+Figure 2: Comparison of the supervised fine-tuning of the single- versus multi-task GOAL on four new problems.
+Figure 3: Unsupervised fine-tuning of GOAL to eight diverse new tasks. We report the best results over 10 runs both for training from scratch and fine-tuning.
+4.3 Ablation Study
+Mixed-attention blocks.
+We compare our mixed-attention blocks with alternative modules which also aim at mixing the edge information with the attention scores, namely MatNet [kwon_matrix_2022] and G2G [henderson_transformers_2023].
+More precisely we focus on the ATSP and train three versions of GOAL with the three different variants of mixed-attention.
+For MatNet, we keep the original hyper-parameters used for solving the ATSP.
+Fig 4 (left) presents the models performance (optimality gap on a test set of 128 instances) at different points of the training.
+We see that using our mixed-attention blocks leads to a significant gain, on the performance at convergence (MatNet) or on the speed of convergence (G2G).
+Note that these ablations are performed on a smaller training dataset than in the previous experiments (100k instances), hence the larger optimality gap.
+Multi-type transformer.
+The proposed multi-type transformer architecture is meant to deal with tasks that are naturally represented by multipartite graphs with heterogeneous nodes or edges, typically scheduling problems where jobs and machine nodes play different roles.
+To evaluate its effect on the performance, we compare it to that obtained with the more direct alternative which ignores node and edge heterogeneity: we simply concatenate the features of the different types of nodes (resp.
+edges), and use dummy “padding” values whenever a feature is undefined for a given node (resp.
+edge), so that the original single-type model can be used.
+We implemented this alternative for the Job Shop Scheduling Problem and present in Fig 4 (middle) a comparison between the training performance of the single-type and multi-type architectures.
+Clearly, the multi-type architecture converges faster and to a better performance: after 200 epochs of training on 100K instances, it reaches a test optimality gap of 4.6% while the single-type model achieves only 8.9%.
+Sparsification using a codebook.
+Fig 4 (right) experimentally confirms the importance, in the fine-tuning phase (supervised mode), of sparsifying the input adapter with a shared codebook.
+The results are given for the PCTSP, but similar results are available in Fig 6 (Appendix) for the other problems.
+The shared codebook results in a more stable finetuning (over 10 runs) and to a better optimality gap.
+While we do not observe a significant effect on the performance on the training tasks, we believe that constraining the rank of the codebook tends to prevent its vectors from specializing for individual tasks seen at training, which would be of limited use when fine-tuning to new tasks.
+Figure 4: (left) Performance with different variants of mixed-attention. (middle) Performance of single- vs multi-type GOAL. (right) Fine-tuning of GOAL with and without the codebook.
+5 Discussion and Conclusion
+In this paper, we propose a generalist neural CO model (GOAL) and demonstrate the feasibility of learning a shared backbone model to solve a variety CO problems, spanning routing, scheduling, packing and classic graph problems.
+We have shown the benefits of GOAL as (i) a single-task specialized model, providing state-of-the-art results on seven CO problems; (ii) a multi-task model, able to solve different problems with the same backbone; and (iii) a pretrained model, which is efficiently fine-tuned to new problems.
+The proposed multi-task pre-training relies on imitation learning and therefore the availability of good-quality solutions.
+While this is not an intrinsic limitation of the GOAL model, we believe that imitation learning is an effective training paradigm for learning multi-task models.
+This is indeed demonstrated by previous works which relied on sequence-modeling training tasks for control [sun_smart_2023, reed_generalist_2022].
+This is particularly advantageous for CO as we can leverage high-quality solutions to classical NP-hard problems, provided by traditional solvers, which are based on decades of OR research.
+In contrast, for a new potentially more original problem, we show that GOAL can be effectively finetuned without any supervision.
+Our approach is limited to problems where solution feasibility is easy to ensure during the construction, which is a common requirement of constructive heuristics.
+Finally the quadratic complexity of the underlying transformer architecture makes the approach too costly for problem sizes beyond a thousand nodes.
+Exploring more efficient forms of attention as basis for our mixed-attention blocks would be a promising direction for better scaling.
+Appendix A GOAL Hyperparameters
+The GOAL backbone consists of 9 layers with an embedding size of 128, a feed-forward dimension of 512, and ReZero normalization [bachlechner_rezero_2021].
+The mixed-attention blocks have 8 heads.
+Low dimensions for nodes and edges are 8 and 4, respectively.
+Appendix B Supervised fine-tuning of GOAL
+In section 4.2 we describe the process of fine-tuning our multitask model on eight unseen tasks in unsupervised mode.
+Although fine-tuning without labeled data is useful for transferring to new tasks, it comes at a cost - it is quite expensive.
+Our experiments show that fine-tuning to the new problems with unsupervised learning takes, on average, a few hours.
+On the other hand, we demonstrate that GOAL can be effectively fine-tuned using a very small amount of labeled data, in a short time.
+Figure 5 presents the performance of fine-tuning GOAL on the same eight problems used in unsupervised tuning.
+Here, we use only 128 solved (but not necessarily optimal) instances and tune the model with just one training step per instance (more precisely, one training step per tail-subproblem of each instance).
+In contrast to unsupervised tuning, this method is extremely fast; fine-tuning takes only a few minutes.
+Figure 5: Supervised fine-tuning on eight new problems. We report the best results over 10 runs for both training from scratch and fine-tuning.
+Appendix C Impact of codebook on fine-tuning
+In 4.3 We discussed the importance of sparsifying the input adapter with a shared codebook.
+Although this feature does not impact training performance, it is very important in the fine-tuning phase.
+By adding a codebook to the model, fine-tuning becomes more stable and achieves better performance.
+In the main text, we presented a figure comparing fine-tuning for the PCTSP with models that have and do not have a codebook.
+Figure 6 illustrates the impact of the codebook on fine-tuning the multitask model across all eight problems we used for transfer.
+In these experiments, we used light-trained models with and without a codebook.
+The models are trained on 100K instances, for 200 epochs.
+For each experiment, we perform 10 inference runs and report the mean with a confidence interval of 100.
+Figure 6: Ablation study of codebook
+Appendix D Generalization to new instance distributions
 Table 2 presents the generalization to instances up to 10 times bigger than the training ones, for a subset of the training tasks.
-We see an excellent generalization performance on ATSP and CVRP.
-For the JSSP, the generalization is not as strong but still very good on most distributions.
-Note that scaling to much larger instances is not a focus of this paper.
-Recent approaches which do not rely on quadratic attention blocks such as [39] would be more appropriate, although highly specialized to some problems.
-We believe that the more expensive mixed-attention blocks are crucial for the model’s capacity to handle diverse tasks.
-We also tested finetuning GOAL to other distribution shifts, namely to the Euclidian distance for TSP and asymmetric distances for CVRP (while GOAL was trained on ATSP and Euclidian CVRP).
-Results are presented in Appendix B and demonstrate a strong performance.
-Train. distribution Generalization
+Our approach takes a bit more time than some baselines, but note that scaling to much larger instances is not a focus of this paper.
+One possible way to avoid the quadratic complexity of attention is by using task-specific heuristics, such as the KNN heuristic for routing problems (proposed by [drakulic_bq-nco_2023], to reduce the size of the input instance and the overall complexity.
+Another approach could be using methods that do not rely on quadratic attention blocks, such as those proposed by [luo_self-improved_2024], though this method tends to be highly specialized for specific problems.
+Our model demonstrates excellent generalization performance in terms of the optimality gap for problems with various edge features, such as ATSP (with an asymmetric distance matrix) and JSSP (with two binary matrices with precedence constraints and job-machine relations).
+Train. distr. Generalization
 ATSP100 ATSP200 ATSP500 ATSP1000
 MatNet greedy 1.57% 1s 97.21% 12s - -
-BQ-NCO greedy 1.63% 1s 1.55% 9s 4.29% 26s 8.09% 1m
-GOAL multi-task greedy 0.45% 5s 1.54% 30s 1.48% 3m 1.89% 1h
+BQ-NCO greedy 1.63% 1s 1.90% 9s 4.29% 26s 8.09% 1m
+GOAL multi-task greedy 0.50% 5s 1.84% 53s 1.77% 3m 2.37% 28m
 CVRP 100 CVRP200 CVRP500 CVRP1000
@@
 BQ-NCO greedy 2.79% 1s 2.81% 9s 3.64% 55s 5.88% 2m
-GOAL multi-task greedy 2.85% 5s 2.80% 12s 3.63% 48s 6.63% 3m
-JSSP10x10 JSSP15x15 JSSP20x15 JSSP20x20
-COMPASS 4.70% 23m 8.00% 38m 10.40% 1h -
-GOAL multi-task greedy 2.65% 6m 5.90% 18m 17.40% 32m 10.80% 59m
+GOAL multi-task greedy 3.06% 5s 3.35% 30s 3.96% 2m 7.63% 3m
 Table 2: Generalization performance on larger instances. Lower is better.
-4.2 Fine-Tuning to New Tasks
-In this section, we want to evaluate the transfer capability of the trained GOAL backbone to unseen tasks.
-We consider four new CO problems: the Traveling Repairman Problem (TRP), the Open Vehicle Routing Problem (OVRP), the Split Delivery Vehicle Routing Problem (SDVRP) and the Maximum Independent Set (MIS) problem (detailed description in Appendix A).
-For each problem, our first attempt was to define new task-specific input and output adapters and train them either with or without back-propagation through the backbone.
-It turned out that in that case, the fine-tuning process is very unstable and has a strong dependency on the parameter initialization (i.e.
-the random seed).
-Next, we initialized the input adapters with the ones of the closest training problems: TSP for TRP, CVRP for OVRP and SDVRP and MVC for MIS.
-The output adapters were still trained from scratch.
-In this case the fine-tuning is very effective.
-In Fig 2, we see that the finetuned model reaches very good optimality gaps (at most $5\%$) much faster than a single-task GOAL model trained from scratch.
-Our few-shot fine-tuning process relies on only 64 instances and lasts less than 1 minute.
-We use the same loss and optimizer as for the training but with a smaller initial learning of 0.0001.
-These results assume that the whole model is fine-tuned, including the backbone.
-When the latter is frozen, we observe a slight degradation of the results with no significant speedup.
-From these experiments, we hypothesize that the input adapters are the difficult part of the transfer.
-To test this hypothesis, we considered one of the training tasks, discarded its input and/or output adapters and tried to re-learn them from scratch, with the trained backbone.
-This led to similar observations: re-learning the output adapter works, while re-learning the input adapter is very unstable.
-We conclude that transfer in GOAL is successful for problems that share the same instance structure as one of the training problems and thus can benefit from a pre-trained input adapter.
-Note that the new problems can have arbitrary objective and constraints (as is the case with the four tested new tasks).
-Appendix D further discusses input adapters and their impact on the latent representations.
-Figure 2: Fine tuning GOAL to new tasks.
-4.3 Ablation Study
-Figure 3: Comparison of the model performance when using different variants of mixed-attention blocks for solving the Asymmetric Traveling Salesman problem
-Figure 4: Comparison between single and multi-type architectures for the Job Shop Scheduling Problem
-Mixed-attention blocks.
-To evaluate the effectiveness of our mixed-attention blocks, we replace them by alternative modules which also aim at mixing the edge information with the attention scores, namely MatNet [34] and G2G [20].
-More precisely we focus on the ATSP and train three versions of GOAL with the three different blocks.
-For MatNet, we keep the original hyper-parameters used for solving the ATSP.
-Figure 4 presents the models performance (optimality gap on a test set of 128 instances) at different points of the training.
-We see that using our mixed-attention blocks leads to a significant gain, on the performance at convergence (MatNet) or on the speed of convergence (G2G).
-Note that these ablations are performed on a smaller training dataset than in the previous experiments (100k instances), hence the wider opt gap.
-Multi-type transformer.
-The proposed multi-typed transformer architecture is meant to deal with tasks that are naturally represented by multipartite graphs with heterogeneous nodes or edges, typically scheduling problems where jobs and machine nodes play different roles.
-To evaluate its effect on the performance, we compare it to that obtained with the more direct alternative which ignores node heterogeneity: we simply concatenate the features of the different types of nodes (or edges) and use dummy values whenever a feature is undefined, so that the original single-type model can be used.
-We implemented this alternative for the Job Shop Scheduling Problem and present in Fig 4 a comparison between the training performance of the single-type and multi-type architectures.
-Clearly the multi-type architecture converges faster and to a better performance: after 200 epochs of training on 100K instances, it reaches a test optimality gap of 4.6% while the single-type model achieves only 8.9%.
-5 Discussion and Conclusion
-In this paper, we propose a generic neural CO model (GOAL) and demonstrate the feasibility of learning a shared backbone model to solve a variety CO problems, spanning routing, scheduling, packing and classic graph problems.
-We have shown the benefits of GOAL as (i) a single-task specialized model, providing state-of-the-art results on six CO problems; (ii) a multi-task model, able to solve several problems with the same backbone; and (iii) a pretrained model, to be efficiently fine-tuned for new problems.
-To the best of our knowledge, this is the first neural approach that handles such a variety of problems with a single backbone model.
-We believe this is a key step in the direction of a foundation model for CO, where the diversity of problem representations (as opposed to the uniformity of textual or visual data) is still a major hurdle.
-Our experiments have exclusively been conducted with Imitation Learning, which assumes the availability of good-quality solutions for at least a certain size of instances.
-While this is not an intrinsic limitation of the GOAL model – the underlying MDPs being defined independently of the training process – we believe that Imitation Learning is an effective training paradigm for learning multi-task models.
-This is indeed demonstrated by previous works which relied on sequence-modeling training tasks for control [51, 47].
-Leveraging solutions to classical NP-hard problems, provided by traditional solvers, which are based on decades of OR research, is particularly advantageous for pretraining a model.
-Such a model can later be adapted to more exotic problems with minimal supervision data, or even better, with Reinforcement Learning.
-Indeed using Reinforcement Learning to fine-tune GOAL to new problems, for which high-quality solvers may not be available, constitutes a promising direction for future work.
-Appendix A Description of the considered CO problems
-In this Section, we describe the CO problems that are solved by GOAL as well as the instance generation process.
-We organize the problems into three categories: Routing, Scheduling and Packing and Graph Problems.
-We also provide an overview Table of the problems and used solvers for each category.
-A.1 Routing Problems
-TSP TRP OP
-Node features - - Price
-Edge features Distance between nodes
-Instance features - - Maximum tour length
-Objective $min$(tour length) $min$(sum of waiting times) $max$(collected prices)
-Oracle Concorde(LKH) LKH EA4OP
-Action Select a node
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-CVRP/SDCVRP CVRPTW OCVRP
-Node features Demands Demands Time window starts Time window ends Service times Demands
-Edge features Distance between nodes
-Instance features Vehicle capacity
-Objective Minimize total tour length
-Oracle Concorde/HGS HGS LKH
-Action Select a next node, subroute is done
-[Uncaptioned image] [Uncaptioned image]
-Table 3: Overview of the routing problems solved by GOAL
-A.1.1 Traveling Salesman Problem
+Appendix E Description of the considered CO problems
+In this Section, we describe the CO problems (in alphabetical order) that are solved by GOAL as well as the instance generation process.
+GOAL solves problems from various categories: Routing, Scheduling, Packing, Graph, and Location Problems.
+For each problem, we provide all the details of how it is represented in our model, including a list of instance, node and edge features, the action chosen by the model, and the objective.
+We also describe how we generate a solution trajectory for problems where we use single cross-entropy loss for behavioral cloning.
+E.1 Capacitated Vehicle Routing Problem and Split-Delivery Capacitated Vehicle Routing Problem
+The Capacitated Vehicle Routing Problem (CVRP) is a standard routing problem.
+The goal of CVRP is to optimize the delivery routes of a fleet of vehicles to a set of customers while considering each vehicle’s capacity.
+The objective is to minimize the overall distance traveled by the vehicles while respecting the capacity constraints and ensuring that all the customer demands are met.
+Split-Delivery CVRP (SDCVRP) is a variant of the CVRP that allows a single customer’s demand to be fulfilled by multiple vehicles.
+In our experiments we use Euclidean CVRP for training GOAL and SDCVRP for fine-tuning.
+Data generation.
+Node locations are randomly sampled from the unit square, while node demands are randomly generated integers from the interval [1, 10].
+Following [kool_attention_2019], the vehicle capacity depends on the problem size: capacity of 50 for a problem size of 100, 80 for 200, 100 for 500, and capacity of 250 for problems with 1000 nodes.
+Instance features: Remaining vehicle capacity
+Node features: Origin/destination tokens, demand
+Edge features: Distance between nodes
+Oracle: LKH
+Objective: Minimize traveled distance
+Action: Select a node, with direct edge or via-depot
+Trajectory: Nodes ordered by a solution subtour. Subtours are ordered by their final
+remaining capacity in descending order
+E.2 Capacitated Vehicle Routing Problem with Time Windows
+The Capacitated Vehicle Routing Problem with Time Windows (CVRPTW) is an extension of the CVRP.
+In addition to demand constraints, each customer also has a specified time window during which they can be visited, along with the necessary service time required for serving that customer.
+The objective of the CVRPTW is to find routes for the vehicles that satisfy the capacity limits, deliver goods within the specified time windows, and minimize the total travel and waiting time or the number of vehicles used.
+In our experiments, we consider the standard version of the Euclidean problem where total traveling and serving time should be minimized and we use it for training GOAL.
+Data generation.
+For the time windows and service times, we use a similar procedure to [falkner_learning_2020], where the locations are sampled sampled from the unit square.
+We ensure the feasibility of the resulting instances by sampling time windows using a more refined procedure.
+Windows start time values are sampled from the range (0.01, 0.25).
+It’s important to note that start times do not directly impact problem feasibility, as vehicles can wait at the node until the service starts.
+On the other hand, the choice of the visiting end time has an impact on feasibility.
+An instance is unfeasible if the window end time of a customer is less than the travel time from the depot plus the service time.
+To avoid this, end times are sampled from the interval [max(window start time, travel time from depot) + 2 $\times$ service time, max(window start time, travel time from depot) + 4 $\times$ service time].
+Instance features: Remaining vehicle capacity, current time
+Node features: Origin/destination tokens, demand, start time window, end time window,
+service time
+Edge features: Distance between nodes
+Oracle: HGS
+Objective: Minimize traveled time
+Action: Select a node, with direct edge or via-depot
+Trajectory: Nodes ordered by a solution subtour. Subtours are ordered by their final
+remaining capacity in descending order
+E.3 Job Shop Scheduling Problem
+The Job Shop Scheduling Problem (JSSP) is a classic scheduling problem.
+We are given $N$ jobs and $M$ machines.
+Each job consists of a sequence of operations (they can called tasks) that need to be processed in a specific order (implying precedence constraints).
+Each operation has a given processing time.
+Each machine can only process one operation at a time and when an operation is started it cannot be interrupted (no preemption).
+A solution is a schedule, defined by the start time of each operation such that a specific criterion, such as makespan, total completion time, or total tardiness is minimized.
+Data generation.
+We follow the standard procedure for generating JSSP benchmark datasets: execution times are randomly sampled from $U(1,100)$.
+Instance features: None
+Node features: Job execution time, job availability time, machine availability time
+Edge features: Precedence constraint, task-machine dependency
+Oracle: ORTools
+Objective: Minimize makespan
+Action: Select a task
+Trajectory: Tasks ordered by minimal finishing time
+E.4 Knapsack Problem
+The Knapsack Problem (KP) is a well-known combinatorial optimization problem that involves selecting a subset of items from a given set.
+Each item had a weight and a value, and the goal is to maximize the cumulated value of selected items, without exceeding a given cumulated weight threshold, known as the knapsack’s (weight) capacity.
+Data generation.
+Following the same procedure as in [kwon_pomo_2020], node weights and values are randomly sampled from the unit interval, and knapsack capacity is set to 25.
+Instance features: Remaining knapsack capacity
+Node features: Value, weight
+Edge features: None
+Oracle: ORTools
+Objective: Maximizing packed values
+Action: Select an item
+Trajectory: There is no order
+E.5 Maximal Independent Set Problem
+In the Maximal Independent Set (MIS), we are given an undirected graph $G=(V,E)$ and the goal is to find a subset of nodes $S\subset V$ of maximal size $|S|$ such that each pair of nodes in $S$ is not connected by an edge.
+Simple speaking, a Maximal Independent Set (MIS) in a graph is a subset of vertices that is independent (no two vertices in the set are adjacent) and is not a subset of any larger independent set.
+In our experiments, we fine-tune GOAL to MIS problem.
+Data generation.
+For this problem, we follow a similar procedure as in [khalil_learning_2017]. We generate random graphs by using Erdős–Rényi model with the critical edge probability $p$ sampled from [0.05, 0.15].
+Instance features: None
+Node features: None
+Edge features: Adjacency value (0-1)
+Oracle: FastWVC
+Objective: Maximize number of selected nodes
+Action: Select an node
+Trajectory: There is no order
+E.6 Maximum Coverage Location Problem
+In the Maximum Coverage Location Problem (MCLP), we are given a set of locations with associated weights and several facilities with associated radius of coverage.
+The goal is to find the optimal placement of the facilities, on a subset of the locations, to maximize the number of covered locations.
+There is also a variant in which nodes have associated weights; in this case, the objective is to maximize the sum of the weights of the covered locations.
+In our experiments, we fine-tune our multitask model to MCLP.
+Data generation.
+Node locations are randomly sampled from the unit square, the number of facilities is fixed to 10 and their radius of coverage to 0.1.
+Instance features: Number of facilities, radius of coverage
+Node features: None
+Edge features: Distance between nodes
+Oracle: PuLP [^2] https://github.com/coin-or/pulp
+Objective: Maximize the number of covered locations
+Action: Select an facility node
+Trajectory: There is no order
+E.7 Minimum Vertex Cover Problem
+The goal of the Minimum Vertex Cover Problem (MVC) is to find the smallest set of vertices in a graph, such that each edge is incident to at least one of these vertices.
+In the weighted version of this problem (MWVC), each vertex has an associated weight, and the goal is to select vertices with the minimum cumulated sum of weights while satisfying the vertex cover condition.
+Data generation and problem features are the same as in MIS.
+We use MVC to train our model.
+E.8 Open Shop Scheduling Problem
+The Open Shop Scheduling Problem (OSSP) is a scheduling problem in which a set of jobs (with a given set of operations) must be processed on a set of machines, but unlike job shop scheduling, there is no predefined order.
+Each operation must be processed, with no overlapping between operations and machines at the same time unit.
+The goal is typically to minimize the makespan, or the total time required to complete all jobs.
+Data generation and problem features are the same as in JSSP, except OSSP does not have precedence constraints.
+E.9 Open Vehicle Routing Problem
+Open Vehicle Routing Problem (OCVRP) is a variant of CVRP where an unbounded fleet of vehicles is available at the depot, each serving a set of customers.
+In contrast with other CVRP variants, vehicles do not need to return to the depot at the end of their service.
+This makes OCVRP similar to the classical CVRP problem, with a different objective function.
+In our experiments, we handle the Euclidean version of the OCVRP and fine-tune GOAL to solve it.
+Data generation.
+Node locations are randomly sampled from the unit square, while node demands are randomly generated integers from the interval [1, 10].
+As in CVRP, the vehicle capacity is set to 50 for a problem size of 100.
+Instance features: Remaining vehicle capacity
+Node features: Origin/destination tokens, demand
+Edge features: Distance between nodes
+Oracle: LKH
+Objective: Minimize traveled distance
+Action: Select a node, with direct edge or via-depot
+Trajectory: Nodes ordered by a solution subtour. Subtours are ordered by their final
+remaining capacity in descending order
+E.10 Orienteering Problem
+The Orienteering Problem (OP) is a combinatorial optimization problem that involves finding the optimal route to maximize the collected prizes while visiting a set of locations and satisfying a maximal distance constraint.
+Data generation.
+We generated data following previous work [kool_attention_2019] and  [xin_multi-decoder_2021]: node locations are uniformly sampled from the unit square and prizes are proportional to the distance to the depot.
+We followed the same procedure line in  [kool_attention_2019], this variant of the problem is called $OP_{distance}$.
+The maximal tour length constraint is set to 4.
+We use Euclidean OP for training GOAL.
+Instance features: Remaining tour length
+Node features: Origin/destination tokens, prize
+Edge features: Distance between nodes
+Oracle: A4OP
+Objective: Maximize collected prizes
+Action: Select a node (including coming back to the depot)
+Trajectory: Ordered nodes by a solution tour
+E.11 Price Collection Traveling Salesman Problem
+The Prize-Collecting Traveling Salesman Problem (PCTSP) is a variant of the classic Traveling Salesman Problem.
+In the PCTSP, each node has an associated prize and penalty, and the goal is to minimize the traveled distance while satisfying a minimum total prize constraint.
+Unlike the traditional TSP, where all cities must be visited, the PCTSP allows for skipping some nodes; however, unvisited nodes incur a penalty, creating a trade-off between collecting prizes and minimizing travel distance.
+The final objective is to minimize the sum of the traveled distance and the sum of the penalties for unvisited nodes.
+Data generation.
+We generated data following previous work [xin_multi-decoder_2021]: node locations are uniformly sampled from the unit square, while prizes are randomly sampled from $U[0,num-nodes]$ and penalties are sampled from $U[0,num-nodes]$.
+Both features are normalized by multiplying by $4/num-nodes$ and $12/num-nodes$, respectively.
+Instance features: Remaining prize constraint
+Node features: Origin/destination tokens, prize, penalty
+Edge features: Distance between nodes
+Oracle: ORTools
+Objective: Minimize the sum of traveled distances and penalties of univisited nodes
+Action: Select a node (including coming back to the depot)
+Trajectory: Ordered nodes by a solution tour
+E.12 Sequential Ordering Problem
+The Sequential Ordering Problem (SOP) is a combinatorial optimization problem where a set of tasks or nodes must be visited in a specific order while minimizing the total travel cost or distance.
+This problem can be modeled as the Traveling Salesman Problem, with precedence constraints.
+The objective is to find the shortest possible route that satisfies these precedence conditions.
+In our experiments, we fine-tune our multitask model to SOP.
+Data generation.
+The values of the cost matrix are integers randomly generated from a uniform distribution $U[1,1000]$ with zero costs between the same task.
+To fulfill feasibility constraints, we randomly order the tasks, and for the $i$-th task, we randomly add $k$ precedent jobs, where $k$ is sampled from $U[0,\lfloor i/2\rfloor]$.
+Instance features: None
+Node features: Origin token
+Edge features: Distance between nodes, precedence constraint
+Oracle: LKH
+Objective: Minimize cost
+Action: Select a node
+Trajectory: Ordered nodes by a solution tour
+E.13 Traveling Repairman Problem
+The Traveling Repairman Problem (TRP) is a combinatorial optimization problem where a repairman must determine the most efficient route to visit a set of locations, performing repairs or services at each site.
+In the Traveling Repairman Problem, a set of locations is given with the time needed to travel between any pair of locations.
+Each location has a task to be performed by the repairman.
+The objective of the problem is to find the route that minimizes the sum of the delays for reaching each point.
+In our experiments, we fine-tune GOAL to Euclidean TRP.
+Data generation.
+Node locations are uniformly sampled from the unit square.
+Instance features: None
+Node features: Origin token
+Edge features: Distance between nodes
+Oracle: LKH
+Objective: Minimize waiting time of all nodes
+Action: Select a node
+Trajectory: Ordered nodes by a solution tour
+E.14 Traveling Salesman Problem
 The Traveling Salesman Problem (TSP) is a classic combinatorial optimization problem.
@@
 A TSP instance can be represented either by node coordinates or by a distance matrix.
-The majority of NCO models address TSP in the coordinate setting, yet the matrix setting holds greater significance for practical applications, as it applies beyond the Euclidean context, where the matrix can represent time, energy or cost, rather than distance, and need not satisfy the symmetry property of a distance matrix.
+The majority of NCO models address TSP in the coordinate setting, yet the matrix setting holds greater significance for practical applications, as it applies beyond the Euclidean context, where the matrix can represent time, energy, or cost, rather than distance, and need not satisfy the symmetry property of a distance matrix.
+In our work, we use the Asymmetric TSP (where the distances satisfy the triangular inequality property) to train our multi-task model."
 Data generation
-For the Euclidian case, we generate datasets as described in [30]: points are uniformly sampled from the unit square, and the corresponding distance matrix $A$ is computed.
-For non-Euclidean problems (ATSP) we generate data as described in [34], where distance matrices are generated randomly, but post-processed to satisfy the triangular inequality.
-A.1.2 Capacitated Vehicle Routing Problem
-The Capacitated Vehicle Routing Problem (CVRP) is another standard routing problem.
-The goal of CVRP is to optimize the delivery routes of a fleet of vehicles to a set of customers, while considering each vehicle’s capacity.
-The objective is to minimize the overall distance traveled by the vehicles, while respecting the capacity constraints and ensuring that all the customer demands are met.
-Data generation.
-The matrix is sampled in the same way as for TSP, while node demands are randomly generated integers from the interval [1, 10].
-Following [30], the vehicle capacity depends on the problem size: capacity of 50 for a problem size of 100, 80 for 200, 100 for 500, and capacity of 250 for problems with 1000 nodes.
-A.1.3 CVRP with Time Windows
-The CVRP with Time Windows (CVRPTW) is an extension of the CVRP.
-In addition to demand constraints, each customer also has a specified time window during which they can be visited, along with the necessary service time required for serving that customer.
-The objective of the CVRPTW is to find routes for the vehicles that satisfy the capacity limits, deliver goods within the specified time windows, and minimize the total travel and waiting time or the number of vehicles used.
-In our model, we consider the standard version of the problem where total traveling and serving time should be minimized.
-Data generation.
-For the time windows and service times, we use a similar procedure to [17], where the travel times between nodes are sampled in the range [0, $\sqrt{2}\pm 0.1$].
-We ensure the feasibility of the resulting instances by sampling time windows using a more refined procedure.
-Windows start time values are sampled from the range (0.01, 0.25).
-It’s important to note that windows start time do not directly impact problem feasibility, as vehicles can wait at the node until the service starts.
-On the other hand, the choice of the windows end time has an impact on feasibility.
-An instance is unfeasible if the window end time of a customer is less than the travel time from the depot plus the service time.
-To avoid this, windows end time are sampled from the interval [max(window start time, travel time from depot) + 2 $\times$ service time, max(window start time, travel time from depot) + 4 $\times$ service time].
-A.1.4 Split Delivery VRP
-The Split Delivery VRP (SDVRP) is similar to the CVRP with the difference that a customer can be visited by several vehicles, i.e. its demand is split between multiple vehicles.
-Data generation.
-We use the same data as for the CVRP (Sec A.1.2).
-A.1.5 Open Vehicle Routing Problem
-This is a variant of CVRP where an unbounded fleet of vehicles is available at the depot, each serving a set of customers, and they do not need to return to the depot at the end of their service.
-Data generation.
-We use the same data as for the CVRP (Sec A.1.2).
-A.1.6 Orienteering Problem
-The Orienteering Problem (OP) is a combinatorial optimization problem that involves finding the optimal route to maximize the collected prizes while visiting a set of locations and satisfying a maximal distance constraint.
-Data generation.
-Again we follow previous work [30]: node locations are sampled in the same way as for the TSP and CVRP and prizes are randomly sampled integers from the interval [1, 10].
-The maximal distance constraint is set to 4.
-A.1.7 Traveling Repairman Problem
-In the travelling repairman problem (TRP), as in TSP, a set of locations is given as well as the time needed to travel between any pair of locations.
-Each location has a task to be performed by the repairman.
-The objective of the problem is to find the route that minimizes the sum of waiting times for all tasks.
-Data generation.
-We use the same data as for the TSP (Sec A.1.1)
-A.2 Packing and Graph Problems
-KP MVC MIS
-Node features Price and weight -
-Edge features - Adjacency matrix
-Instance features Knapsack capacity -
-Objective $max$(collected prices) $max$(vertex coverage) $min$(independent set)
-Oracle ORTools FastWVC
-Action Select a node
-Table 4: Overview of the packing and graph problems solved by GOAL
-A.2.1 Knapsack Problem
-The Knapsack Problem (KP) is a well-known combinatorial optimization problem that involves selecting a subset of items from a given set.
-Each item had a weight and a value, and the goal is to maximize the cumulated value of selected items, but without exceeding a given cumulated weight threshold, known as the knapsack’s (weight) capacity.
-Data generation.
-Following the same procedure as in [33], node weights and values are randomly sampled from the unit interval and knapsack capacity is set to 25 for all problem sizes.
-A.2.2 Minimum (Weighted) Vertex Cover Problem
-The goal of the Minimum Vertex Cover Problem (MVC) is to find the smallest set of vertices in a graph, such that each edge is incident to at least one of these vertices.
-In the weighted version of this problem (MWVC), each vertex has an associated weight, and the goal is to select vertices with the minimum cumulated sum of weights while satisfying the vertex cover condition.
-Data generation.
-For this problem, we follow a similar procedure as in [27]. We generate random graphs by using Erdős–Rényi model with the critical edge probability $p$ sampled from [0.05, 0.15].
-A.2.3 Maximum Independent Set
-In the Maximum Independent Set (MIS), we are given an undirected graph $G=(V,E)$ and the goal is to find a subset of nodes $S\subset V$ of maximal size $|S|$ such that each pair of nodes in $S$ is not connected by an edge.
-Data generation.
-We use the same data generation process as for the Minimum (Weighted) Vertex Cover Problem.
-A.3 Scheduling Problems
-UMSP JSSP
-Node features Machine: Availability time Task: Execution time, Availability time Machine: Availability time
-Edge features Job-machine: Execution times Task-task: precedence constraints, $ij=1$ if tasks $i$ and $j$ belong to the same job Task-machine: $ij=1$ if task $i$ must be executed on machine $j$
-Objective Minimize a makespan
-Oracle HiGHS ORTools
-Action Select a machine for a random task Select a task
-[Uncaptioned image] [Uncaptioned image]
-Table 5: Overview of the scheduling problems solved by GOAL
-A.3.1 Job Shop Scheduling Problem
-The Job Shop Scheduling Problem (JSSP) is a classic scheduling problem.
-We are given $N$ jobs and $M$ machines.
-Each job consists of a sequence of operations that need to be processed in a specific order (implying precedence constraints).
-Each operation has a given processing time.
-Each machine can only process one operation at a time and when an operation is started it cannot be interrupted (no preemption).
-A solution is a schedule, defined by the start time of each operation such that the makespan, i.e.
-the maximum completion time of the operations, is minimized.
-Data generation.
-We follow the standard procedure for generating JSSP benchmark datasets: execution times are randomly sampled from $U(1,10)$.
-A.3.2 Unrelated Machine Scheduling Problem
-In the Unrelated Machine Scheduling problem (UMSP), the goal is to schedule $N$ jobs on $M$ machines such that the makespan is minimized.
-Each job $i$ has a processing time $p_{ij}$ on each machine $j$, which is arbitrary (hence the term unrelated).
-There are no precedence constraints and the schedules are non-preemptive.
-The solution to the problem is essentially a partition of the jobs on the machines.
-[21] shows that this problem is NP-hard.
+For the Euclidean case, we generate datasets as described in [kool_attention_2019]: points are uniformly sampled from the unit square, and the corresponding distance matrix $A$ is computed.
+For non-Euclidean problems (ATSP) we generate data as described in [kwon_matrix_2022], where distance matrices are generated randomly, but post-processed to satisfy the triangular inequality.
+Instance features: None
+Node features: Origin/destination tokens
+Edge features: Distance between nodes
+Oracle: Concorde (for Euclidean) and LKH (for non-Euclidean)
+Objective: Minimize tour length
+Action: Select a node
+Trajectory: Ordered nodes by a solution tour
+E.15 Unrelated Machine Scheduling Problem
+The Unrelated Machine Scheduling Problem (UMSP) is an optimization problem in operations research where a set of jobs must be assigned to a set of machines, each with different processing times for each job, to minimize a specific criterion, such as makespan, total completion time, or total tardiness.
+Unlike related machine scheduling, where machines have similar characteristics and processing times are proportional, in UMSP, the performance and efficiency of machines vary, making it a more complex and challenging problem to solve.
+[horowitz_exact_1976] shows that this problem is NP-hard.
+We use UMSP for multitask training of our GOAL model.
 Data generation.
 For UMSP, the processing times are randomly generated from $U(1,100)$.
-Appendix B Finetuning of GOAL on Various Distribution Shifts
-We consider the following distribution shifts: TSP with Euclidian distances and CVRP with asymmetric distances (while GOAL was trained on ATSP and Euclidian CVRP).
-Fig 5 compares finetuning GOAL on these new distributions, versus training a new single-task model from scratch.
-We observe that finetuning converges faster and to a better optimality gap.
-Figure 5: Tuning GOAL to various distribution shifts
-Appendix C GOAL Hyperparameters
-The GOAL backbone consists of 9 layers with an embedding size of 128, a feed-forward dimension of 512, and ReZero normalization [3]. The mixed-attention blocks have 8 heads.
-Appendix D Analysis of the latent representations
-To investigate the interaction between the different CO problems, we compare the latent representations of the different tasks and their evolution through the layers of the trained GOAL model.
-Since these are high-dimensional vectors, we use the standard t-SNE technique [53] to visualize the representations in the 2D space.
-More precisely, in Fig 6 we plot the t-SNE projection (with perplexity $150$) of the node embeddings after the first, fourth, and last layer of the model, for each of the training tasks (top) and with the addition of the transfer tasks (bottom).
-For each task, we plot 20 full instances of size 100 (110 for scheduling problems).
-We observe that the representations seem clustered per problem in the different layers.
-We also see clear overlaps between similar problems (CVRP-SDCVRP-OVRP in green shades and TSP-TRP in blue shades) in layers 1 and 4 while they become separated in the last layer.
-These visualizations indicate that the problem-specific input adapters seem to map the instances to distinct areas of the embedding space.
-This might explain the strong dependency between the input adapters and the backbone which makes learning a new input adapter for a new task often challenging.
-Figure 6: Visualization of the evolution of the node latent representations through the GOAL model for the training tasks (top) and with the addition of the transfer tasks (bottom)
+Instance features: None
+Node features: Machine availability times
+Edge features: Execution times
+Oracle: ORTools
+Objective: Minimize makespan
+Action: Select a machine to randomly selected task
+Trajectory: Ordered tasks by execution time

### v3

@@
-GOAL: A Generalist Combinatorial Optimization Agent Learning
+GOAL: A Generalist Combinatorial Optimization Agent Learner
 Abstract
@@
 Any variation of a problem requires adjustment of its model and re-training from scratch.
-In this paper, we propose GOAL (for Generalist combinatorial Optimization Agent Learning), a generalist model capable of efficiently solving multiple COPs and which can be fine-tuned to solve new COPs.
+In this paper, we propose GOAL (for Generalist combinatorial Optimization Agent Learner), a generalist model capable of efficiently solving multiple COPs and which can be fine-tuned to solve new COPs.
 GOAL consists of a single backbone plus light-weight problem-specific adapters for input and output processing.
@@
 Finally we showcase the strong transfer learning capacity of GOAL by fine-tuning it on several new problems.
-Our code is available at https://github.com/naver/goal-co/.
+Our code is available at this url.
 1 Introduction
 Combinatorial Optimization (CO) problems are extensively studied in the Operations Research community, because they are both challenging in terms of theoretical complexity and central in many applications such as transportation, logistics and finance.
-Traditional CO methods include exact solvers (typically based on integer programming [korte_combinatorial_2018]) and approximation algorithms [vazirani_approximation_2003] which provide optimality or approximation guarantees, respectively, but do not scale; and heuristic algorithms [boussaid_survey_2013] which produce good-quality solutions faster but without guarantees.
+Traditional CO methods include exact solvers, typically based on integer programming [korte_combinatorial_2018], and approximation algorithms [vazirani_approximation_2003] which provide optimality or approximation guarantees, respectively, but do not scale; and heuristic algorithms [boussaid_survey_2013] which produce good-quality solutions faster but without guarantees.
 Effective heuristics heavily rely on expert knowledge and are generally problem-specific.
 A more recent and promising family of methods for CO is based on Machine Learning, either to learn the hyper-parameters of classic CO solvers, or to directly learn heuristics in an end-to-end fashion [bengio_machine_2021, cappart_combinatorial_2021].
-In particular constructive neural heuristics have been successful at solving a variety of CO problems, including routing [bello_neural_2017, nazari_reinforcement_2018, kool_attention_2019, kwon_pomo_2020], scheduling [zhang_learning_2020, kwon_matrix_2022] and graph problems [khalil_learning_2017, qiu_dimes_2022].
+In particular constructive neural heuristics have been successful at solving a variety of CO problems, including routing [bello_neural_2017, nazari_reinforcement_2018, kool_attention_2018, kwon_pomo_2020], scheduling [zhang_learning_2020, kwon_matrix_2021] and graph problems [khalil_learning_2017, qiu_dimes_2022].
 However, in most cases, a separate model is specialized and trained for each problem, and sometimes even for each distribution of instances within a problem, especially for varying instance sizes.
@@
 Graphs, unlike images or text contain little redundancy or intrinsic semantics.
-In this paper, we propose a new model, named GOAL (for Generalist combinatorial Optimization Agent Learning), which consists of a single backbone and light-weight problem-specific adapters for input and output processing.
+In this paper, we propose a new model, named GOAL (for Generalist combinatorial Optimization Agent Learner), which consists of a single backbone and light-weight problem-specific adapters for input and output processing.
 To handle the large discrepancy between CO problems, we propose three complementary strategies.
@@
 We further provide an ablation study to evaluate the effectiveness of a shared codebook, multi-type transformer architecture and mixed-attention blocks.
-Notably the GOAL architecture gives state-of-the-art results on 7 out of our 8 training tasks when used in as single-task model.
+Notably the GOAL architecture gives state-of-the-art results on 7 out of our 8 training tasks when used as single-task model.
 In summary, our contributions are the following:
@@
 Our work falls into the category of end-to-end constructive heuristics for combinatorial optimization.
-Constructive neural heuristics have been successfully applied to a wide range of COPs, including many variants of routing problems [kool_attention_2019, xin_multi-decoder_2021, kim_sym-nco_2022, kool_deep_2021, fu_generalize_2021, sun_difusco_2023], scheduling problems [zhang_learning_2020, kwon_matrix_2022], as well as classic COPs such as the Knapsack Problem [kwon_pomo_2020] and the Maximum Independent Set [khalil_learning_2017, sun_difusco_2023].
-Another successful category of neural CO methods are improvement heuristics, which focus on iteratively enhancing a given complete solution, typically by using a neural network to guide the selection of local search operators [chen_learning_2019, hottung_neural_2019, ma_learning_2021, wu_learning_2020].
-The seminal work of [khalil_learning_2017] introduces graph embeddings which do work for a variety of CO problems but with limited performance.
+Constructive neural heuristics have been successfully applied to a wide range of COPs, including many variants of routing problems [kool_attention_2018, xin_multi-decoder_2021, kim_sym-nco_2022, kool_deep_2022, fu_generalize_2021, sun_difusco_2023], scheduling problems [zhang_learning_2020, kwon_matrix_2021], as well as classic COPs such as the Knapsack Problem [kwon_pomo_2020] and the Maximum Independent Set [khalil_learning_2017, sun_difusco_2023].
+Another successful category of neural CO methods are improvement heuristics, which focus on iteratively enhancing a given complete solution, typically by using a neural network to guide the selection of local search operators [chen_learning_2019, hottung_neural_2020, ma_learning_2021, wu_learning_2022].
+The seminal work of khalil_learning_2017 introduces graph embeddings which do work for a variety of CO problems but with limited performance.
 Current state-of-the-art models are tailored for certain problem structures (e.g.
-euclidian routing problems, where nodes are represented by their coordinates plus some extra node or graph features such as demand or vehicle capacity), and successfully handled by transformer-based architectures [kool_attention_2019].
-For non-euclidian routing problems, [kwon_matrix_2022] and [drakulic_bq-nco_2023] propose adapted architectures to deal with the distance matrix, but they are specialized for such problems and do not apply to those with node-only features.
+euclidian routing problems, where nodes are represented by their coordinates plus some extra node or graph features such as demand or vehicle capacity), and successfully handled by transformer-based architectures [kool_attention_2018].
+For non-euclidian routing problems, kwon_matrix_2021 and drakulic_bq-nco_2023 propose adapted architectures to deal with the distance matrix, but they are specialized for such problems and do not apply to those with node-only features.
 In GOAL, we propose a novel mixed-attention modules which can flexibly deal with both node and edge features, incurring no overhead when one or the other is missing.
@@
 While the first seminal models for neural CO had a dramatic drop in performance when tested on out-of-training distribution instances [joshi_learning_2022, manchanda_generalization_2022], several recent approaches show a greatly improved generalization performance, especially to larger instances [fu_generalize_2021, li_learning_2021, son_meta-sage_2023, pan_h-tsp_2023], to distribution shifts of other instance parameters [bi_learning_2022], or both [manchanda_generalization_2022, bi_learning_2022, zhou_towards_2023, drakulic_bq-nco_2023].
-Concurrently with our work, some papers have started to explore the cross-problem generalization for the family of routing problems.
-Specifically, [lin_cross-problem_2024] propose to pretrain the AM model [kool_attention_2019] on the TSP and to fine-tune either the entire model or some light-weight modules on the Price Collecting TSP and OP.
-Additionally, [zhou_mvmoe_2024, liu_multi-task_2024, berto_routefinder_2024] explore generalization to different variants of the CVRP.
+Concurrently with our work, some papers have started to explore cross-problem generalization for the family of routing problems.
+Specifically, [lin_cross-problem_2024] propose to pretrain the AM model [kool_attention_2018] on the TSP and to fine-tune either the entire model or some light-weight modules on the Price Collecting TSP and OP.
+Additionally, zhou_mvmoe_2024, liu_multi-task_2024, berto_routefinder_2024 explore generalization to different variants of the CVRP.
 They build on the idea of zero-shot compositionality [ruis_independent_2021] by training a model on CVRPs with some combinations of constraints from a predefined set (e.g.
 time windows and duration limit constraints) and show that the model generalizes to some unseen combinations of the same constraints.
-Although these approaches tackle multiple problems, it is important to note that the considered Euclidian routing problems remain very similar in structure and indeed many of those problems were solved by the seminal Attention Model of [kool_attention_2019] with little adaptations of its base model.
+wang_efficient_2023 propose a bandit based approach to alternate between problems during training, so as to focus on those which have the best return.
+Although these approaches tackle multiple problems, it is important to note that the considered Euclidian routing problems remain very similar in structure and indeed many of those problems were solved by the seminal Attention Model of [kool_attention_2018] with little adaptations of its base model.
 Our paper goes further and considers multi-task training and transfer learning of a new model on a variety of COPs with different structures.
@@
 Although they treat various tasks, with potentially diverse state and action spaces, these works still consider image observations and exploit this common structure.
-[reed_generalist_2022] proposes GATO, a general-purpose agent obtained by training a large transformer-based model on hundreds of tasks spanning vision, language and control.
+reed_generalist_2022 propose GATO, a general-purpose agent obtained by training a large transformer-based model on hundreds of tasks spanning vision, language and control.
 It demonstrates the potential of a single network (albeit with 1.2B parameters) to effectively tackle a wide variety of tasks.
@@
 Figure 1: Left: Architecture of GOAL for single-type problems; the green components are task-specific, while the others belong to the backbone.
-The codebook is a shared $L{\times}D$ (resp.
-$\bar{L}{\times}\bar{D}$) matrix applied to the node (resp.
+The codebook is a shared ×$\ell{\times}D$ (resp.
+¯×$\bar{\ell}{\times}\bar{D}$) matrix (with $\ell{\ll}D,\bar{\ell}{\ll}\bar{D}$) applied to the node (resp.
 edge) representations produced by the input adapters.
-Right: Architecture of GOAL for multi-type problems (with here 2 node types T1,T2); “self” and “cross” are identical multi-head mixed attention blocks (same parameters per layer for all types and tasks); “FF” for the two types are also identical Feed Forward blocks per layer.
-For the edges, all the type combinations are shown (here T11,T12,T21,T22) but each task can define its own.
+Right: Architecture of GOAL for multi-type problems (with here 2 node types T1,T2); it uses multiple mixed attention blocks (“MMA”), some in self- and some in cross-attention mode, but they all share the same parameters (per layer); so do the two Feed Forward blocks (“FF”).
+For the edges, all the type combinations are shown (here T11,T12,T21,T22) but those which have no meaning in a task are simply omitted.
 The same set of parameters is used for the two architectures.
@@
 Thus, solving the problem instance can be formulated as a sequential decision procedure which iteratively selects the construction steps so as to reach not only a feasible but an optimal solution.
-That in turn can be modeled as a Markov Decision Problem (MDP) associated with the problem instance.
-In this paper, we use the BQ-MDPs introduced by [drakulic_bq-nco_2023] which are defined for the widespread class of tail-recursive problems  [bertsekas_dynamic_vol1_2012].
-In BQ-MDPs, the actions are the construction steps, typically the selection of a node, while the states are instances.
+That in turn can be modeled as a Markov Decision Problem (MDP) associated with the problem.
+In this paper, we use the BQ-MDPs introduced by drakulic_bq-nco_2023 which are defined for “tail-recursive” problems, i.e.
+the widespread class of problems that are solvable by dynamic programming [bertsekas_reinforcement_2019].
+In BQ-MDPs, the actions are the construction steps, typically the selection of a node, while the states are simply instances.
+This differs from many peer NCO models where each state is a pair of an instance and a partial solution.
+The tail recursion property precisely ensures that the two alternatives lead to “bisimilar” MDPs, which have the same optimal policies, as shown in drakulic_bq-nco_2023, but the BQ-MDPs’ states more efficiently capture the relevant information to determine these policies.
 3.2 Multi-Task CO Problems
@@
 $\bar{F}_{t}$ of the $N{\times}N$ edges).
-While some tasks may have no edge feature ($\bar{F}_{t}{=}0$), we always assume that at least one node feature exists, which distinguishes the nodes by a randomly assigned value between 00 and $1$, acting as node identifier, so we always have $F_{t}{\geq}1$.
+While some tasks may have no edge feature ($\bar{F}_{t}{=}0$), we always assume that at least one node feature exists, which distinguishes the nodes by a randomly assigned value between 00 and $1$, similar to the random features used in Graph Neural Networks [sato_random_2021], so we always have $F_{t}{\geq}1$.
 Both $F_{t}$ and $\bar{F}_{t}$, as well as the meaning of the corresponding features, depend on task $t$.
@@
 The layers themselves, on the other hand, are only aware of the shared embedding dimensions $D,\bar{D}$.
-They are self-attention transformer layers [vaswani_attention_2017] where vanilla attention is replaced by multi-head mixed attention, similar in intent to [kwon_matrix_2022, henderson_transformers_2023] but with a different design.
+They are self-attention transformer layers [vaswani_attention_2017] where vanilla attention is replaced by multi-head mixed attention, similar in intent to [kwon_matrix_2021, henderson_transformers_2023] but with a different design.
 Our mixed attention block, in the general case where $M$ nodes attend on $N$ nodes ($M{=}N$ only in self-attention), is an instance of generalized attention [andreoli_convolution_2019].
@@
 $Q_{n}^{(h)}$ $=Q_{n}\mathbf{W}_{Q}^{(h)},$ $K_{m}^{(h)}$ $=K_{m}\mathbf{W}_{K}^{(h)},$ $V_{m}^{(h)}$ $=V_{m}\mathbf{W}_{V}^{(h)}.$
-Then the key and query vectors are used to compute an $M{\times}N$ score matrix $S$, which is transformed into an attention matrix (by adding the mask and applying softmax) multiplied by the reprojected value vectors to produce the output representation $r$ of the $N$ input query nodes as an $N{\times}D$ matrix:
+Then the key and query vectors are used to compute an $M{\times}N$ score matrix $S$, which is transformed into an attention matrix (by adding the mask and applying softmax) multiplied by the reprojected value vectors to produce the output representation $r$ of the $N$ input query nodes as a matrix with $N$ rows:
 $r$ $=\sum_{h}\underbrace{\textrm{softmax}_{\textrm{col}}(S^{(h)}{+}\mathcal{M})^{\top}}_{\textrm{attention matrix}}V^{(h)}\mathbf{W}_{O}^{(h)\top}.$
@@
 Our mixed attention allows to incorporate arbitrary relational information (edge features) into the core of the attention mechanism.
-It differs from positional encoding, which attempts to distill a priori the edge information (limited to sequential ordering) into the node representations.
+It differs from positional encoding, which attempts to instill a priori the edge information (limited to sequential ordering) into the node representations.
 In the GOAL backbone, all the node (resp.
@@
 Whenever available, we use such heuristics.
-Input and output adapters are linear projections, of the task-specific instance features to the backbone’s input embedding (dim.
-$D$) and from the backbone’s output embedding to task-specific action space.
-However, when $D$ is large and the number of training tasks is small, there is a risk that the embeddings produced by each task end up occupying complementary subspaces of the embedding space throughout the layers, which would defeat the purpose of effectively sharing parameters across tasks.
+Input and output adapters are linear projections, from the task-specific instance features to the backbone’s input embeddings and from the backbone’s output embeddings to the task-specific action space.
+However, when the embedding dimension is large and the number of training tasks is small, there is a risk that the embeddings produced by each task end up occupying complementary subspaces of the embedding space throughout the layers, which would defeat the purpose of effectively sharing parameters across tasks.
 To prevent this, we simply force the input linear projection to be low rank.
-Concretely, node (resp.
-edge) features are first projected into a “small” dimension space, forcing the representations to share dimensions across tasks, then this low dimension representation is plunged into the “large” $D$-dimensional embedding space using a shared codebook (see Fig.
+Thus, node (resp.
+edge) features are first mapped into a low dimension $\ell{\ll}D$ (resp.
+¯$\bar{\ell}{\ll}\bar{D}$), through a task-specific linear projection (the input adapter proper), forcing the representations to share dimensions across tasks.
+Then this “small” representation is plunged into a “large” $D$- (resp.
+$\bar{D}$)-dimensional embedding, through a common linear projection, called a codebook, shared by all the tasks (see Fig.
 1 left).
@@
 This is illustrated in Fig.
-1 right.
+1 right in the case of two types.
+For $n$ types, the architecture would be similar but involves $n^{2}$ mixed-attention blocks, where each type attends to itself and every other type, or a selection of relevant types defined by the task.
 Therefore, in our scheduling example, the model will contain separate self-attention blocks, for operation nodes and machine nodes respectively, and cross-attention blocks, for operation-machine edges and vice-versa.
@@
 This is precisely what allows us to train a unique model to solve arbitrary single- and multi-type problems.
+3.6 Training Procedure
+We train GOAL on a set of eight various CO problems (described in Sec 4).
+For each training problem, we define the corresponding BQ-MDP, the linear input and output adapters to match the dimensions of the input features and the construction actions.
+We train the model by imitation of high-quality solutions, provided by specialized solvers.
+At each step, we sample a problem and create a batch of instances by sampling suffix subproblems from the training instances, similarly to [drakulic_bq-nco_2023].
+The forward and backward processes involve the input and output adapters, as well the backbone parameters, where the mixed-attention and feed-forward blocks are duplicated within each layer if the sampled task is multi-typed.
 4 Experiments
@@
 The training instances are of size 100 nodes, or 10$\times$10 for JSSP and 100$\times$20 for UMSP (resulting in both cases in 100 operation nodes).
-The description of the problems and the associated data generation processes is provided in Appendix E.
-Training process.
+The description of the problems and the associated data generation processes is provided in Appendix F.
+Training details.
 GOAL is trained by alternating batches in random order from each of the training tasks, using a single-class cross-entropy loss for ATSP, CVRP, CVRPTW, OP, JSSP and UMSP and a multi-class cross-entropy loss for KP and MVC.
-Similarly to [drakulic_bq-nco_2023], we create a batch by sampling suffix subproblems for each instance.
-We use AdamW optimizer [loshchilov_decoupled_2019], with an initial learning rate of 0.0005 and a decay rate of 0.97 every 10th epoch.
+We use AdamW optimizer [loshchilov_decoupled_2018], with an initial learning rate of 0.0005 and a decay rate of 0.97 every 10th epoch.
 We trained the model on 8 Nvidia V100 GPU servers, with batch size of 256 for 7 days, completing ca.
@@
 Baselines.
-We consider the following constructive neural models as baselines: AM [kool_attention_2019], MDAM [xin_multi-decoder_2021], POMO [kwon_pomo_2020], Sym-NCO [kim_sym-nco_2022], BQ-NCO [drakulic_bq-nco_2023], MVMoE [zhou_mvmoe_2024], MatNet [kwon_matrix_2022], S2V-DQN [khalil_learning_2017] and COMPASS [chalumeau_combinatorial_2023].
-All models except MVMoE are specialized and trained in single-task mode.
-They are often combined at test time with various enhancers, such as beam search [drakulic_bq-nco_2023], simulation-guided beam search  [choo_simulation-guided_2022], instance augmentation [kwon_pomo_2020], or active search [hottung_efficient_2022], but for fair comparison, all the results we report are obtained by a pure greedy application of the trained policy at test time.
+We consider the following constructive neural models as baselines: AM [kool_attention_2018], MDAM [xin_multi-decoder_2021], POMO [kwon_pomo_2020], Sym-NCO [kim_sym-nco_2022], BQ-NCO [drakulic_bq-nco_2023], MVMoE [zhou_mvmoe_2024], MatNet [kwon_matrix_2021], S2V-DQN [khalil_learning_2017], RouteFinder [berto_routefinder_2024], COMPASS [chalumeau_combinatorial_2023] and Gumbeldore [pirnay_self-improvement_2024].
+All models except MVMoE and RouteFinder are specialized and trained in single-task mode.
+They are often combined at test time with various enhancers, such as beam search [drakulic_bq-nco_2023], simulation-guided beam search [choo_simulation-guided_2022], MCTS [xing_graph_2020], instance augmentation [kwon_pomo_2020], or active search [hottung_efficient_2021], but for fair comparison, all the results we report are obtained by a pure greedy application of the trained policy at test time.
 4.1 Performance on the Training Tasks
@@
 More broadly, they demonstrate that multi-task learning can be a successful paradigm for neural combinatorial optimization.
-In addition, in Appendix D, we provide experimental results demonstrating GOAL’s excellent zero-shot generalization on instances up to 10 times larger than the training instances.
+In addition, in Appendix E, we provide experimental results demonstrating GOAL’s excellent zero-shot generalization on instances up to 10 times larger than the training instances.
 We can note in particular the strong generalization for ATSP which seems much more challenging for the state-of-the-art baselines.
@@
 ATSP100 CVRP100 CVRPTW100 OP100
-opt. gap time opt. gap time opt. gap time opt. gap time
+gap time gap time gap time gap time
 Oracle solver 0.00% 29s 0.00% 12m 0.00% 10m 0.00% 1.1m
@@
 MVMoE/4E - 1.65% 1s 4.90% 1s -
+RouteFinder-TE - 1.50% 1s 3.19% 1s -
 MatNet greedy 0.93% 1s - - -
-GOAL single-task greedy 0.21% 10s 2.31% 10s 2.79% 10s -0.35% 3s
-GOAL multi-task greedy 0.50% 10s 3.06% 10s 2.86% 11s 0.28% 3s
+GOAL single-task greedy 0.30% 10s 2.34% 10s 2.61% 10s -0.04% 3s
+GOAL multi-task greedy 0.91% 10s 3.16% 10s 3.82% 10s 0.43% 3s
 KP100 MVC100 UMSP100x20 JSSP10x10
-opt. gap time opt. gap time opt. gap time opt. gap time
-Oracle solver 0.00% 1s 0.00% 2m 0.00% 2m 0.00% 12m
+gap time gap time gap time gap time
+Oracle solver 0.00% 1s 0.00% 2m 0.00% 2m 0.00% 47s
 POMO no aug 0.19% - - -
@@
 S2V-DQN - 0.97% 2s - -
-COMPASS - - - 4.70% 23m
-GOAL single-task greedy 0.10% 3s 0.26% 3s 1.30% 7s 2.70% 15s
-GOAL multi-task greedy 0.14% 2s 0.37% 9s 2.01% 7s 3.74% 15s
-Table 1: Performance on 128 test instances from the training tasks. Lower is better.
+COMPASS - - - 4.70% 3h
+Gumbeldore greedy - - - 3.17% 9s
+GOAL single-task greedy 0.10% 3s 0.23% 3s 2.82% 7s 2.73% 15s
+GOAL multi-task greedy 0.12% 3s 0.37% 3s 3.84% 7s 4.13% 15s
+Table 1: Performance on 128 test instances from the training tasks. Each gap is with respect to the oracle and often corresponds to the optimality gap. Lower is better.
 4.2 Fine-Tuning to New Tasks
 In this section, we want to evaluate the transfer capability of the trained GOAL backbone to unseen tasks.
-We consider eight diverse new COPs: the Traveling Repairman Problem (TRP), the Prize Collecting TSP (PCTSP), the Open CVRP (OCVRP), the Split Delivery CVRP (SDCVRP), the Sequential Ordering Problem (SOP), the Maximum Covering Location Problem (MCLP), the Maximal Independent Set (MIS) problem and the Open Shop Scheduling Problem (OSSP) (detailed descriptions in Appendix E).
+We consider eight diverse new COPs: the Traveling Repairman Problem (TRP), the Prize Collecting TSP (PCTSP), the Open CVRP (OCVRP), the Split Delivery CVRP (SDCVRP), the Sequential Ordering Problem (SOP), the Maximum Covering Location Problem (MCLP), the Maximal Independent Set (MIS) problem and the Open Shop Scheduling Problem (OSSP) (detailed descriptions in Appendix F).
 GOAL can be fine-tuned to specific problems in two modes.
@@
 This is similar to the self-labeling strategy proposed by corsini_self-labeling_2024 for the JSSP.
-One could also use more sophisticated experts such as generic local search heuristics or smarter sampling pirnay_self-improvement_2024 to accelerate the finetuning.
+One could also use more sophisticated experts such as generic local search heuristics or smarter sampling [pirnay_self-improvement_2024] to accelerate the finetuning.
 Performance results are given in Fig 2, which shows that unsupervised fine-tuning systematically outperforms unsupervised training of the whole model from scratch at a comparable processing time budget.
@@
 CVRP model for OCVRP or JSSP model for OSSP), hinting to the synergistic effect of the multi-task training.
-Figure 2: Comparison of the supervised fine-tuning of the single- versus multi-task GOAL on four new problems.
 Figure 2: Unsupervised fine-tuning of GOAL to eight diverse new tasks. We report the best results over 10 runs both for training from scratch and fine-tuning.
@@
 Mixed-attention blocks.
-We compare our mixed-attention blocks with alternative modules which also aim at mixing the edge information with the attention scores, namely MatNet [kwon_matrix_2022] and G2G [henderson_transformers_2023].
+We compare our mixed-attention blocks with alternative modules which also aim at mixing the edge information with the attention scores, namely MatNet [kwon_matrix_2021] and G2G [henderson_transformers_2023].
 More precisely we focus on the ATSP and train three versions of GOAL with the three different variants of mixed-attention.
@@
 Multi-type transformer.
-The proposed multi-type transformer architecture is meant to deal with tasks that are naturally represented by multipartite graphs with heterogeneous nodes or edges, typically scheduling problems where jobs and machine nodes play different roles.
-To evaluate its effect on the performance, we compare it to that obtained with the more direct alternative which ignores node and edge heterogeneity: we simply concatenate the features of the different types of nodes (resp.
-edges), and use dummy “padding” values whenever a feature is undefined for a given node (resp.
-edge), so that the original single-type model can be used.
-We implemented this alternative for the Job Shop Scheduling Problem and present in Fig 4 (middle) a comparison between the training performance of the single-type and multi-type architectures.
+Our multi-type architecture is meant to deal with tasks that are naturally represented by multipartite graphs with heterogeneous nodes or edges, typically scheduling problems where job and machine nodes play different roles.
+To evaluate its effect on the performance, we compare it to that obtained with the more direct alternative which ignores node and edge heterogeneity: we simply gather in a single tensor the node (edge) representations of the different types, produced by the input adapter, and use the original single-type model.
+Note that because of the linearity of the input adapters, these representations can be obtained either by defining a specific linear transform per type, or equivalently by defining a single linear transform on the direct sum of the feature spaces of the different types.
+We implemented the former for the Job Shop Scheduling Problem and present in Fig 3 (middle) a comparison between the training performance of the single-type and multi-type architectures.
 Clearly, the multi-type architecture converges faster and to a better performance: after 200 epochs of training on 100K instances, it reaches a test optimality gap of 4.6% while the single-type model achieves only 8.9%.
+Finally, note that the multi-type model offers more flexibility for fine-tuning: while, in order to be task agnostic at training, parameter sharing between the type specific attention blocks of a layer must be enforced, that constraint can be relaxed at fine-tuning to better fit the task at hand.
 Sparsification using a codebook.
@@
 Figure 3: (left) Performance with different variants of mixed-attention. (middle) Performance of single- vs multi-type GOAL. (right) Fine-tuning of GOAL with and without the codebook.
+Figure 4: Comparison of the (supervised) fine-tuning of the single- versus multi-task GOAL on four new problems.
 5 Discussion and Conclusion
@@
 This is indeed demonstrated by previous works which relied on sequence-modeling training tasks for control [sun_smart_2023, reed_generalist_2022].
-This is particularly advantageous for CO as we can leverage high-quality solutions to classical NP-hard problems, provided by traditional solvers, which are based on decades of OR research.
+It is particularly advantageous for CO as we can leverage high-quality solutions to classical NP-hard problems, provided by traditional solvers, which are based on decades of OR research.
 In contrast, for a new potentially more original problem, we show that GOAL can be effectively finetuned without any supervision.
@@
 Exploring more efficient forms of attention as basis for our mixed-attention blocks would be a promising direction for better scaling.
-Appendix A GOAL Hyperparameters
-The GOAL backbone consists of 9 layers with an embedding size of 128, a feed-forward dimension of 512, and ReZero normalization [bachlechner_rezero_2021].
+Appendix A GOAL Hyperparameters and size
+The GOAL backbone consists of $L{=}9$ layers with an embedding size of $D{=}\bar{D}{=}128$ (the same for both nodes and edges), a feed-forward dimension of 512, and ReZero normalization [bachlechner_rezero_2021].
 The mixed-attention blocks have 8 heads.
-Low dimensions for nodes and edges are 8 and 4, respectively.
+The task specific input adapters produce representations of low dimension $\ell{=}8$ for nodes and ¯$\bar{\ell}{=}4$ for edges, which are mapped to embeddings by a shared codebook of dimension ×$\ell{\times}D$ for nodes and ¯×$\bar{\ell}{\times}\bar{D}$ for edges, respectively.
+GOAL backbone has 2.1M parameters and adapters are very light, with only a few thousand parameters per task.
+For the illustration, RouteFinder [berto_routefinder_2024] with POMO has 1.3M, MVMoE [zhou_mvmoe_2024] has 3.7M (the same model is used in RouteFinder with MVMoE), and RouteFinder with the Transformer Encoder 1.7M.
 Appendix B Supervised fine-tuning of GOAL
@@
 Figure 6: Ablation study of codebook
-Appendix D Generalization to new instance distributions
+Appendix D Multi-type mixed attention layers for UMSP and JSSP
+As an illustration of our proposed multi-type transformer architecture, we present in Figure 7 the inputs to the mixed-attention blocks for the Unrelated Machine Scheduling (UMSP) and the Job Shop Scheduling Problems (JSSP).
+Note that in the case of UMSP, since there is no edge information between jobs and between machines, the self-attention blocks (left-most and right-most attention blocks) degenerate to vanilla attention.
+The cross-attention blocks, on the other hand, are mixed, and they are fed the same edge information, holding the job execution times on each machine (transposed from each other).
+In JSSP, the situation is similar, but ops have two kinds of relations between themselves: precedence, and sharing the same job.
+Hence the self-attention block on ops is mixed.
+(a) the Unrelated Machine Scheduling Problem (UMSP)
+(b) the Job-shop Scheduling Problem (JSSP)
+Figure 7: Illustration of the multi-type model on scheduling tasks in GOAL. The "MA" blocks use vanilla multi-head attention while the "MMA" use our multi-head mixed-attention.
+Appendix E Generalization to new instance distributions
 Table 2 presents the generalization to instances up to 10 times bigger than the training ones, for a subset of the training tasks.
@@
 Our model demonstrates excellent generalization performance in terms of the optimality gap for problems with various edge features, such as ATSP (with an asymmetric distance matrix) and JSSP (with two binary matrices with precedence constraints and job-machine relations).
-Train. distr. Generalization
-ATSP100 ATSP200 ATSP500 ATSP1000
-MatNet greedy 1.57% 1s 97.21% 12s - -
-BQ-NCO greedy 1.63% 1s 1.90% 9s 4.29% 26s 8.09% 1m
-GOAL multi-task greedy 0.50% 5s 1.84% 53s 1.77% 3m 2.37% 28m
-CVRP 100 CVRP200 CVRP500 CVRP1000
-AM greedy 7.11% 1s 10.16% 1s 17.2% 1s 72.95% 2.1s
-MDAM greedy 4.84% 1s 7.54% 6s 12.94% 24s 32.25% 1m
-POMO no aug 1.21% 1s 5.94% 6s 31.36% 32s 284.58% 1m
-Sym-NCO greedy 3.33% 1s 8.42% 3s 37.58% 54s 481.53% 9m
-BQ-NCO greedy 2.79% 1s 2.81% 9s 3.64% 55s 5.88% 2m
-GOAL multi-task greedy 3.06% 5s 3.35% 30s 3.96% 2m 7.63% 3m
-Table 2: Generalization performance on larger instances. Lower is better.
-Appendix E Description of the considered CO problems
+Train. distrib. Generalization
+100 200 500 1000
+ATSP LKH 0.00% 29s 0.00% 48s 0.00% 2m 0.00% 4m
+MatNet 1.57% 1s 97.21% 12s - -
+BQ-NCO 1.63% 1s 1.90% 9s 4.29% 26s 8.09% 1m
+GOAL 0.45% 5s 1.54% 30s 1.62% 3m 1.96% 5m
+CVRP LKH 0.00% 12m 0.00% 2h 0.00% 5h 0.00% 7h
+AM 7.11% 1s 10.16% 1s 17.2% 1s 72.95% 2s
+MDAM 4.84% 1s 7.54% 6s 12.94% 24s 32.25% 1m
+POMO 1.21% 1s 5.94% 6s 31.36% 32s 284.58% 1m
+Sym-NCO 3.33% 1s 8.42% 3s 37.58% 54s 481.53% 9m
+BQ-NCO 2.79% 1s 2.81% 9s 3.64% 55s 5.88% 2m
+GOAL 3.16% 5s 3.42% 30s 3.90% 2m 7.37% 5m
+CVRPTW HGS (1s) $\dagger$ 108/128 2m 14/128 2m 0/128 2m 0/128 2m
+HGS (1m) 0.00% 10m 0.00% 26m 0.00% 2h 0.00% 2h
+GOAL 3.82% 5s 7.71% 30s 9.49% 2m 6.75% 5m
+OP EA4OP 0.00% 1m 0.00% 3m 0.00% 11m 0.00% 45m
+AM 5.35% 1s 11.05% 1s 22.29% 1s 30.73% 1s
+MDAM 2.88% 16s 5.55% 1s 18.31% 4s 30.05% 15s
+SymNCO 2.03% 2s 6.60% 3s 19.81% 24s 31.79% 1m
+GOAL 0.43% 3s 1.60% 10s 7.45% 17s 15.62% 30s
+KP ORTools 0.00% $<1$s 0.00% $<1$s 0.00% $<1$s 0.00% $<1$s
+POMO 0.19% 1s 0.50% 1s 6.41% 31s 5.34% 1m
+BQ-NCO 0.10% 2s 0.14% 6s 0.74% 7s 0.92% 13s
+GOAL 0.12% 3s 1.63% 6s 2.40% 10s 2.59% 16s
+10x10 10x15 15x15 20x15
+JSSP ORTools 0.00% 47s 0.00% 2m 0.00% 3h 0.00% 80h
+COMPASS $\diamond$ 4.70% 3h - 8.00% 5h 10.40% 8h
+GOAL 4.13% 15s 11.02% 43s 10.74% 45s 22.53% 4m
+Table 2: Generalization performance on larger instances.
+Lower is better.
+All NCO methods perform greedy decoding, except for methods marked with $\diamond$; they do not provide results for greedy decoding.
+Solvers marked with $\dagger$ are unable to solve all instances within the given time budget; we report the number of instances successfully solved.
+Appendix F Description of the considered CO problems
 In this Section, we describe the CO problems (in alphabetical order) that are solved by GOAL as well as the instance generation process.
@@
 We also describe how we generate a solution trajectory for problems where we use single cross-entropy loss for behavioral cloning.
-E.1 Capacitated Vehicle Routing Problem and Split-Delivery Capacitated Vehicle Routing Problem
+F.1 Capacitated Vehicle Routing Problem and Split-Delivery Capacitated Vehicle Routing Problem
 The Capacitated Vehicle Routing Problem (CVRP) is a standard routing problem.
@@
 Node locations are randomly sampled from the unit square, while node demands are randomly generated integers from the interval [1, 10].
-Following [kool_attention_2019], the vehicle capacity depends on the problem size: capacity of 50 for a problem size of 100, 80 for 200, 100 for 500, and capacity of 250 for problems with 1000 nodes.
+Following [kool_attention_2018], the vehicle capacity depends on the problem size: capacity of 50 for a problem size of 100, 80 for 200, 100 for 500, and capacity of 250 for problems with 1000 nodes.
 Instance features: Remaining vehicle capacity
@@
 remaining capacity in descending order
-E.2 Capacitated Vehicle Routing Problem with Time Windows
+F.2 Capacitated Vehicle Routing Problem with Time Windows
 The Capacitated Vehicle Routing Problem with Time Windows (CVRPTW) is an extension of the CVRP.
@@
 remaining capacity in descending order
-E.3 Job Shop Scheduling Problem
+F.3 Job Shop Scheduling Problem
 The Job Shop Scheduling Problem (JSSP) is a classic scheduling problem.
 We are given $N$ jobs and $M$ machines.
-Each job consists of a sequence of operations (they can called tasks) that need to be processed in a specific order (implying precedence constraints).
+Each job consists of a sequence of operations that need to be processed in a specific order (implying precedence constraints).
 Each operation has a given processing time.
@@
 Instance features: None
-Node features: Job execution time, job availability time, machine availability time
-Edge features: Precedence constraint, task-machine dependency
+Node features: Operation execution time, job availability time, machine availability time
+Edge features: Precedence constraint, operation-machine dependency
 Oracle: ORTools
 Objective: Minimize makespan
-Action: Select a task
+Action: Select an operation
 Trajectory: Tasks ordered by minimal finishing time
-E.4 Knapsack Problem
+F.4 Knapsack Problem
 The Knapsack Problem (KP) is a well-known combinatorial optimization problem that involves selecting a subset of items from a given set.
@@
 Trajectory: There is no order
-E.5 Maximal Independent Set Problem
+F.5 Maximal Independent Set Problem
 In the Maximal Independent Set (MIS), we are given an undirected graph $G=(V,E)$ and the goal is to find a subset of nodes $S\subset V$ of maximal size $|S|$ such that each pair of nodes in $S$ is not connected by an edge.
@@
 Trajectory: There is no order
-E.6 Maximum Coverage Location Problem
+F.6 Maximum Coverage Location Problem
 In the Maximum Coverage Location Problem (MCLP), we are given a set of locations with associated weights and several facilities with associated radius of coverage.
@@
 Trajectory: There is no order
-E.7 Minimum Vertex Cover Problem
+F.7 Minimum Vertex Cover Problem
 The goal of the Minimum Vertex Cover Problem (MVC) is to find the smallest set of vertices in a graph, such that each edge is incident to at least one of these vertices.
@@
 We use MVC to train our model.
-E.8 Open Shop Scheduling Problem
+F.8 Open Shop Scheduling Problem
 The Open Shop Scheduling Problem (OSSP) is a scheduling problem in which a set of jobs (with a given set of operations) must be processed on a set of machines, but unlike job shop scheduling, there is no predefined order.
@@
 Data generation and problem features are the same as in JSSP, except OSSP does not have precedence constraints.
-E.9 Open Vehicle Routing Problem
+F.9 Open Vehicle Routing Problem
 Open Vehicle Routing Problem (OCVRP) is a variant of CVRP where an unbounded fleet of vehicles is available at the depot, each serving a set of customers.
 In contrast with other CVRP variants, vehicles do not need to return to the depot at the end of their service.
-This makes OCVRP similar to the classical CVRP problem, with a different objective function.
+This makes OCVRP similar to the classical CVRP, as both problems involve a set of locations with demand and vehicle capacity constraints.
+However, the key difference lies in the route structure: in OCVRP, vehicles do not need to return to the depot after completing deliveries, which affects the solution space and cost minimization strategy.
+OCVRP has common applications in logistics, where delivery trucks do not need to return to the depot after completing deliveries.
+Data generation and problem features are the same as for the CVRP.
 In our experiments, we handle the Euclidean version of the OCVRP and fine-tune GOAL to solve it.
-Data generation.
-Node locations are randomly sampled from the unit square, while node demands are randomly generated integers from the interval [1, 10].
-As in CVRP, the vehicle capacity is set to 50 for a problem size of 100.
-Instance features: Remaining vehicle capacity
-Node features: Origin/destination tokens, demand
-Edge features: Distance between nodes
-Oracle: LKH
-Objective: Minimize traveled distance
-Action: Select a node, with direct edge or via-depot
-Trajectory: Nodes ordered by a solution subtour. Subtours are ordered by their final
-remaining capacity in descending order
-E.10 Orienteering Problem
-The Orienteering Problem (OP) is a combinatorial optimization problem that involves finding the optimal route to maximize the collected prizes while visiting a set of locations and satisfying a maximal distance constraint.
-Data generation.
-We generated data following previous work [kool_attention_2019] and  [xin_multi-decoder_2021]: node locations are uniformly sampled from the unit square and prizes are proportional to the distance to the depot.
-We followed the same procedure line in  [kool_attention_2019], this variant of the problem is called $OP_{distance}$.
-The maximal tour length constraint is set to 4.
+F.10 Orienteering Problem
+The Orienteering Problem (OP) is a combinatorial optimization problem that combines aspects of the Traveling Salesman Problem (TSP) and the Knapsack Problem.
+It is commonly used in routing and scheduling applications where the goal is to maximize collected rewards while respecting a constraint on total travel distance or time.
+The OP is defined by a set of locations (nodes), each associated with a reward; a starting location and an ending location (usually the same node); a travel cost matrix between locations; and a maximum allowable travel budget.
+The objective is to visit a subset of locations that maximizes the total collected reward while ensuring the total travel cost does not exceed the budget.
+We generated data following previous work [kool_attention_2018]: node locations are uniformly sampled from the unit square and prizes are proportional to the distance to the depot (in  [kool_attention_2018] this variant of the problem is called $OP_{distance}$).
+The maximal distance constraint is set to 4.
 We use Euclidean OP for training GOAL.
-Instance features: Remaining tour length
+Instance features: Remaining distance
 Node features: Origin/destination tokens, prize
@@
 Trajectory: Ordered nodes by a solution tour
-E.11 Price Collection Traveling Salesman Problem
+F.11 Price Collection Traveling Salesman Problem
 The Prize-Collecting Traveling Salesman Problem (PCTSP) is a variant of the classic Traveling Salesman Problem.
@@
 Trajectory: Ordered nodes by a solution tour
-E.12 Sequential Ordering Problem
+F.12 Sequential Ordering Problem
 The Sequential Ordering Problem (SOP) is a combinatorial optimization problem where a set of tasks or nodes must be visited in a specific order while minimizing the total travel cost or distance.
@@
 Trajectory: Ordered nodes by a solution tour
-E.13 Traveling Repairman Problem
+F.13 Traveling Repairman Problem
 The Traveling Repairman Problem (TRP) is a combinatorial optimization problem where a repairman must determine the most efficient route to visit a set of locations, performing repairs or services at each site.
@@
 Trajectory: Ordered nodes by a solution tour
-E.14 Traveling Salesman Problem
+F.14 Traveling Salesman Problem
 The Traveling Salesman Problem (TSP) is a classic combinatorial optimization problem.
@@
 Data generation
-For the Euclidean case, we generate datasets as described in [kool_attention_2019]: points are uniformly sampled from the unit square, and the corresponding distance matrix $A$ is computed.
-For non-Euclidean problems (ATSP) we generate data as described in [kwon_matrix_2022], where distance matrices are generated randomly, but post-processed to satisfy the triangular inequality.
+For the Euclidean case, we generate datasets as described in [kool_attention_2018]: points are uniformly sampled from the unit square, and the corresponding distance matrix $A$ is computed.
+For non-Euclidean problems (ATSP) we generate data as described in [kwon_matrix_2021], where distance matrices are generated randomly, but post-processed to satisfy the triangular inequality.
 Instance features: None
@@
 Trajectory: Ordered nodes by a solution tour
-E.15 Unrelated Machine Scheduling Problem
+F.15 Unrelated Machine Scheduling Problem
 The Unrelated Machine Scheduling Problem (UMSP) is an optimization problem in operations research where a set of jobs must be assigned to a set of machines, each with different processing times for each job, to minimize a specific criterion, such as makespan, total completion time, or total tardiness.